### v2

# Multi-Preference Lambda-weighted Listwise DPO for Dynamic Preference Alignment

Abstract
While large-scale unsupervised language models (LMs) capture broad world knowledge and reasoning capabilities, steering their behavior toward desired objectives remains challenging due to the lack of explicit supervision. Existing alignment techniques, such as reinforcement learning from human feedback (RLHF), rely on training a reward model and performing reinforcement learning to align with human preferences. However, RLHF is often computationally intensive, unstable, and sensitive to hyperparameters.
To address these limitations, Direct Preference Optimization (DPO) was introduced as a lightweight and stable alternative, enabling direct alignment of language models with pairwise preference data via classification loss. However, DPO and its extensions generally assume a single static preference distribution, limiting flexibility in multi-objective or dynamic alignment settings.
In this paper, we propose a novel framework: Multi-Preference Lambda-weighted Listwise DPO, which extends DPO to incorporate multiple human preference dimensions (e.g., helpfulness, harmlessness, informativeness) and enables dynamic interpolation through a controllable simplex-weighted formulation. Our method supports both listwise preference feedback and flexible alignment across varying user intents without re-training. Empirical and theoretical analysis demonstrates that our method is as effective as traditional DPO on static objectives while offering greater generality and adaptability for real-world deployment.
Introduction
Large language models (LLMs) trained on large-scale text corpora using next-token prediction objectives have demonstrated strong capabilities across a range of natural language understanding and generation tasks. However, these models often exhibit misalignment with human values and intentions, producing responses that may be factually incorrect, verbose, offensive, or otherwise undesirable. This gap arises in part because the standard unsupervised training objective does not explicitly encode human preferences or task-specific constraints.
To improve alignment with human intent, recent research has introduced fine-tuning techniques that incorporate human feedback. Reinforcement learning from human feedback (RLHF) is one prominent approach: it first trains a reward model from pairwise human comparisons, and then optimizes the language model via reinforcement learning to maximize this reward while maintaining similarity to the original policy. While effective, RLHF is known to be unstable, expensive to train, and sensitive to hyperparameter choices. It also requires online sampling and reward estimation, making it cumbersome to deploy at scale.
Direct Preference Optimization (DPO) was proposed as a simpler alternative to RLHF. Rather than relying on reinforcement learning, DPO directly optimizes the model’s output distribution to reflect human preferences using a pairwise classification objective. DPO has been shown to produce high-quality generations with significantly lower training complexity and resource demands. Despite its advantages, the original DPO formulation is limited in its ability to handle diverse or conflicting user preferences, as it is designed for a single fixed reward distribution.
In practical applications, however, alignment often involves multiple criteria—such as helpfulness, harmlessness, factuality, and succinctness—that may vary in importance depending on the use case or user. This motivates a more flexible framework capable of learning from multiple preference dimensions and generalizing across different preference trade-offs.
In this work, we introduce Multi-Preference Lambda-weighted Listwise DPO, a generalization of the DPO framework that incorporates multiple dimensions of preference feedback and enables interpolation over them via a weight vector $\lambda$ drawn from the probability simplex. Our method extends DPO from pairwise to listwise settings and supports alignment under dynamically changing preference configurations. At inference time, users or downstream systems can modulate the alignment behavior by selecting different $\lambda$ vectors without re-training the model.
We present a detailed theoretical derivation of our method starting from the KL-regularized RL objective, and show how it leads to a clean listwise loss that aggregates multiple reward dimensions. We demonstrate the effectiveness of our approach on multi-objective alignment benchmarks, showing that it achieves competitive or superior results compared to conventional DPO, while offering greater flexibility and control.
Related Work
Reinforcement Learning from Human Feedback (RLHF).
RLHF [19, 1] has become the de facto paradigm for aligning large language models (LLMs) with human preferences. This framework typically proceeds in three steps: (1) collect human comparisons over model outputs, (2) train a reward model to mimic these preferences, and (3) fine-tune the base model via reinforcement learning, most commonly using Proximal Policy Optimization (PPO) [22]. While effective in improving response helpfulness and safety, RLHF is complex to implement and computationally intensive, requiring online rollouts, heavy sampling, and sensitive hyperparameter tuning [31]. Moreover, it often suffers from reward hacking and policy collapse [4], making it difficult to scale or adapt in practical applications.
Direct Preference Optimization (DPO).
To address the limitations of RLHF, Direct Preference Optimization (DPO) [20] reparameterizes the reward model such that the optimal policy is analytically defined as a function of preference probabilities. This reformulation allows replacing RL with a simple binary classification objective on preference-labeled data. DPO avoids on-policy rollouts, maintains training stability, and significantly reduces computational cost. Several extensions have since been proposed, such as Quark [27] and LLaMA-DPO [16], which demonstrate competitive or superior performance to PPO-based RLHF in tasks like summarization, question answering, and dialogue generation. However, these approaches are limited to pairwise preferences and assume a single fixed alignment objective.
Listwise Preference Learning.
Traditional learning-to-rank methods distinguish between pairwise [3], pointwise [17], and listwise [25] approaches. Listwise methods are known to better capture global ranking structure, and have recently been adopted for preference modeling in LLMs. UltraFeedback [8] introduces a benchmark dataset where human raters assign rankings to multiple outputs per prompt. Listwise approaches in alignment, such as GLOW [26], use these rankings to learn better reward functions. Our work builds upon these insights but optimizes a listwise loss directly in policy space rather than via supervised reward modeling, enabling end-to-end fine-tuning.
Multi-Objective Alignment.
In many alignment scenarios, objectives such as helpfulness, harmlessness, factuality, and informativeness can conflict. Prior work has addressed this challenge via multi-reward reinforcement learning [12], reward decomposition [13], or manually interpolating between multiple reward functions [24]. However, these methods often require training a separate model for each configuration of preference weights. By contrast, our method incorporates preference weight vectors $\lambda$ during training and samples over the probability simplex, enabling the model to generalize to arbitrary trade-offs without retraining. This mechanism draws inspiration from meta-learning and universal value function approximators [21], but is specifically designed for alignment with human judgments.
Controllable and Steerable LMs.
Beyond RLHF, several techniques aim to control model behavior explicitly. These include prompt-based methods such as zero-shot prompting [2], few-shot prompting [30], and chain-of-thought [23], as well as supervised fine-tuning with control tokens [11] or conditioning variables [29]. Constrained decoding [18] and plug-and-play language models [7] provide post-hoc control. However, these methods typically address narrow or static attributes, lacking the flexibility to express complex human preferences. Our method introduces a theoretically grounded and fine-grained approach for multi-dimensional preference steerability without modifying prompts or decoding strategies.
Preliminaries
We begin by reviewing the foundation of Direct Preference Optimization (DPO) and its connection to KL-regularized reinforcement learning from human feedback (RLHF). Our setting assumes the existence of a base language model $\pi_{\text{ref}}$, a supervised fine-tuned (SFT) model, and a collection of ranked or pairwise human preferences. We aim to train a target policy $\pi_{\theta}$ that better aligns with human preferences while remaining close to $\pi_{\text{ref}}$.
Pairwise Preference Modeling.
Let $x$ denote the prompt, and $y_{w}\succ y_{l}$ represent a human preference between a pair of completions: $y_{w}$ (the preferred answer) and $y_{l}$ (the less preferred one). These preferences are sampled from an underlying distribution $p^{*}(y_{w},y_{l}|x)$, which is typically assumed to be induced by a latent reward function $r(x,y)$. A commonly adopted model is the Bradley-Terry (BT) model [bradley1952rank]:
$p^{*}(y_{w}\succ y_{l}\mid x)=\frac{\exp(r(x,y_{w}))}{\exp(r(x,y_{w}))+\exp(r(x,y_{l}))}$ (1)
Given a dataset of comparisons $\mathcal{D}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$, we can learn the reward function $r(x,y)$ by minimizing the negative log-likelihood under the BT model:
$\mathcal{L}_{\text{BT}}(r)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma(r(x,y_{w})-r(x,y_{l}))\right],$ (2)
where $\sigma(z)=1/(1+\exp(-z))$ is the logistic sigmoid.
KL-Regularized RL Objective.
Once a reward model is trained, the standard RLHF approach fine-tunes the policy $\pi_{\theta}$ to maximize expected rewards while constraining deviation from the reference policy $\pi_{\text{ref}}$ via KL regularization:
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}\left[r(x,y)-\beta\cdot\mathrm{KL}\left(\pi_{\theta}(y|x)\|\pi_{\text{ref}}(y|x)\right)\right],$ (3)
where $\beta>0$ controls the strength of the KL constraint.
This objective encourages the policy to produce high-reward outputs while staying close to $\pi_{\text{ref}}$. However, in practice, it requires sampling, reward evaluation, and reinforcement learning algorithms such as PPO [22] to optimize.
Reparameterization and DPO.
DPO sidesteps explicit reward modeling and reinforcement learning by recognizing that the optimal policy $\pi_{r}$ under Equation 3 can be written as:
$\pi_{r}(y|x)\propto\pi_{\text{ref}}(y|x)\cdot\exp\left(\frac{1}{\beta}r(x,y)\right).$ (4)
Using Equation 4, we reparameterize the reward difference into the log-ratio of policies:
$\beta(r(x,y_{w})-r(x,y_{l}))=\log\frac{\pi_{r}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{r}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}.$ (5)
Replacing $\pi_{r}$ with $\pi_{\theta}$ yields the DPO loss:
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right].$ (6)
This loss directly optimizes the policy parameters $\theta$ using preference comparisons, without sampling or reinforcement learning. It forms the foundation upon which we build our multi-preference and listwise extensions in later sections.
Multi-Preference Lambda-weighted Listwise DPO
Motivated by the limitations of pairwise DPO under a single fixed preference distribution, we introduce a generalized framework that supports listwise preference supervision across multiple human objectives. In real-world settings, human judgments often reflect diverse criteria—such as helpfulness, harmlessness, and conciseness—and capturing these via a unified formulation enables controllable alignment across axes of interest.
Our key insight is that we can reparameterize multi-dimensional listwise preferences using a convex combination of multiple reward signals. Specifically, we define a simplex-weighted listwise loss that extends the DPO framework to dynamic, multi-objective optimization.
Listwise Preference Modeling.
Let $\{y_{1},y_{2},\dots,y_{N}\}$ be a set of candidate completions for a prompt $x$, and let $p^{*}(y_{i}|x)$ be a listwise human preference distribution over these candidates. Analogous to the pairwise DPO reparameterization, we define the listwise policy distribution under $\pi_{\theta}$ as:
$P_{\theta}(y_{i}|x)=\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}},$ (7)
which defines a softmax over reweighted log-ratios, consistent with KL-derived optimal policies.
We then define the listwise DPO objective as a cross-entropy loss between human preferences $p^{*}(y_{i}|x)$ and model predictions:
$\mathcal{L}_{\text{listwise}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[\sum_{i=1}^{N}p^{*}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right].$ (8)
Lambda-weighted Preference Composition.
To enable multi-objective control, we assume access to $m$ distinct preference dimensions, each defined by a distribution $p^{*(k)}(y_{i}|x)$ for objective $k\in\{1,\dots,m\}$. For each prompt, we sample or specify a weight vector $\lambda\in\Delta^{m}$ (the probability simplex), and define the aggregated preference distribution:
$p^{\lambda}(y_{i}|x)=\sum_{k=1}^{m}\lambda_{k}\cdot p^{*(k)}(y_{i}|x).$ (9)
This allows us to interpolate among various alignment goals without retraining.
The full Lambda-weighted Listwise DPO objective becomes:
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right].$ (10)
This expectation can be approximated by sampling $\lambda\sim\text{Unif}(\Delta^{m})$ or conditioning on user-specified preferences.
What does $\lambda$-DPO update do?
To understand the optimization dynamics, consider the gradient of the loss:
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}|x)-P_{\theta}(y_{i}|x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\right].$ (11)
This gradient updates the model to increase likelihood of responses aligned with weighted human preferences, while decreasing likelihood of responses overrepresented in the reference policy. Importantly, when $\lambda$ is fixed (e.g., $\lambda=(1,0,0)$), the objective reduces to a single-objective listwise DPO loss.
Lambda-DPO Pipeline.
The training process proceeds as follows:
  1. 1.
    For each prompt $x$, obtain $N$ candidate completions $\{y_{1},\dots,y_{N}\}\sim\pi_{\text{ref}}(\cdot|x)$.
  2. 2.
    For each alignment dimension $k$, elicit a human listwise preference distribution $p^{*(k)}(y_{i}|x)$.
  3. 3.
    Sample or define $\lambda\in\Delta^{m}$, and compute the mixture preference $p^{\lambda}$.
  4. 4.
    Optimize $\pi_{\theta}$ using the listwise cross-entropy loss between $p^{\lambda}$ and $P_{\theta}$.
In practice, we can reuse datasets like UltraFeedback [8], where multiple preference criteria are annotated. During inference, changing $\lambda$ enables personalized alignment without modifying model parameters.
Theoretical Analysis of Multi-Preference Lambda-weighted Listwise DPO
While Direct Preference Optimization (DPO) offers a robust framework for aligning language models with human preferences via pairwise supervision, it inherently simplifies the preference space by assuming binary feedback between two completions. In this section, we analyze the theoretical implications of this simplification and motivate the use of a listwise and multi-preference extension.
Pairwise DPO and Its Limitations.
Given a prompt $x$ and two outputs $y_{w}$ (winner) and $y_{l}$ (loser), DPO defines a binary preference distribution:
$P^{*}(y_{i}|x)=\begin{cases}1&\text{if }y_{i}=y_{w}\\ 0&\text{if }y_{i}=y_{l}\end{cases}$ (12)
This implicitly constrains the learning objective to satisfy $P^{*}(y_{w}|x)+P^{*}(y_{l}|x)=1$, ignoring the quality of other plausible candidates. Consequently, the optimized policy $\pi_{\theta}$ may fail to capture the full spectrum of user preferences, introducing estimation error.
Listwise Distribution via Softmax.
To mitigate this, we define a listwise target distribution $\pi^{*}$ over all completions $\{y_{1},y_{2},\dots,y_{N}\}$ for prompt $x$, derived from scalar reward scores $r(x,y_{i})$:
$\pi^{*}(y_{i}\mid x)=\frac{\exp(r(x,y_{i}))}{\sum_{j=1}^{N}\exp(r(x,y_{j}))}$ (13)
This distribution captures relative preference over all candidates, allowing for more representative updates and lower variance.
Multi-Preference Supervision.
Real-world alignment involves multiple competing objectives, such as helpfulness, safety, and factuality. Let $p^{*(k)}(y_{i}|x)$ denote the listwise preference under objective $k\in\{1,\dots,m\}$. We define a convex combination of these objectives via a preference vector $\lambda\in\Delta^{m}$:
$p^{\lambda}(y_{i}\mid x)=\sum_{k=1}^{m}\lambda_{k}\cdot p^{*(k)}(y_{i}\mid x)$ (14)
The training objective minimizes the cross-entropy between the aggregated distribution $p^{\lambda}$ and the policy-induced distribution $P_{\theta}$:
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (15)
Gradient Interpretation.
The gradient of this objective is:
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$ (16)
This update rule drives $\pi_{\theta}$ toward the weighted human preference distribution $p^{\lambda}$, allowing alignment to reflect fine-grained and multi-dimensional values.
Variance Reduction via Listwise Supervision.
Using a normalized softmax distribution $\pi^{*}$ derived from reward scores avoids the instability and high variance that arise when optimizing based on hard binary labels. Empirically, listwise training improves sample efficiency and convergence by utilizing the full output ranking structure.
Summary.
In summary, Multi-Preference Lambda-weighted Listwise DPO:
  • Recovers a more informative supervision signal via $\pi^{*}$.
  • Reduces estimation error and variance in gradient updates.
  • Enables alignment over a continuous space of user-defined objectives via $\lambda$.
Experiment Setup
Experimental Objectives and Comparison Targets
The primary goal of our experiments is to evaluate the performance of the proposed $\lambda$-DPO algorithm compared to the standard DPO method. Specifically, we assess their effectiveness across different model architectures and training settings using public preference datasets and standard evaluation benchmarks.
We consider two levels of comparison:
Model comparison: We experiment with two strong open-source language models—LLaMA3-8B and Qwen. These models differ in architecture, pretraining corpora, and tokenizer, providing a robust testbed for evaluating generalizability.
Method comparison: We compare DPO and $\lambda$-DPO. For each method and model pair, we perform independent hyperparameter tuning (e.g., learning rate, batch size, $\beta$) to ensure that comparisons are fair and reflect the best achievable performance under comparable resource constraints.
Training Settings
Inspired by the SimPO baseline setup, we adopt two training pipelines for both models:
  • Base: The model is directly trained on the UltraFeedback dataset using DPO or $\lambda$-DPO. No supervised fine-tuning is applied before preference optimization. This setting isolates the influence of the preference learning method alone.
  • Finetuned-Base: The model is first supervised fine-tuned (SFT) on the UltraChat-200k dataset, which contains approximately 200K high-quality dialogue examples. The resulting SFT model is then optimized using UltraFeedback preferences via DPO or $\lambda$-DPO. This pipeline reflects realistic alignment procedures.
All training is conducted on A100 GPUs, with consistent batch sizes and sequence lengths. For each setting, hyperparameters are tuned via validation performance on reward preference accuracy and KL stability.
Evaluation Benchmarks and Metrics
To comprehensively evaluate model behavior, we use six established benchmarks:
  • MMLU [10]: 57 subject areas with around 15K test questions designed to test academic and professional knowledge.
  • ARC (Challenge) [5]: Grade-school level science questions ($\sim$1.1K) focusing on multi-step reasoning.
  • HellaSwag [28]: 10K commonsense sentence completion tasks with subtle distractors.
  • TruthfulQA [15]: 817 questions designed to expose model hallucination and fact inconsistency.
  • Winograd [14]: 273 examples targeting pronoun coreference and syntactic disambiguation.
  • GSM8K [6]: 8.5K arithmetic reasoning problems requiring multi-step solutions.
We evaluate models along several axes:
  • Win Rate (WR): The proportion of pairwise comparisons where the model’s output is preferred by the reward model.
  • lm-eval-harness Score [9]: Average score given by GPT-4 on multi-turn QA prompts, covering helpfulness, correctness, and language quality.
  • Reward Accuracy (%): The percentage of outputs that match the preferred choice of the reward model.
  • KL Divergence: The average KL divergence between the trained policy and the SFT policy, indicating policy shift magnitude.
  • Run Time: Total training duration under fixed hardware settings.
  • Peak GPU Memory: Maximum memory usage during training.
  • Spearman Correlation: Correlation between output log-probability and output length, used to measure length bias.
Ablation Experiment
To better understand the impact of the $\lambda$ vector in $\lambda$-DPO, we conduct two complementary sets of ablation experiments, each targeting a different aspect of the framework’s flexibility and effectiveness.
The first set of experiments aims to isolate the effect of explicit preference vector interpolation by comparing $\lambda$-DPO models trained with and without multi-dimensional preference aggregation. In this setting, the random $\lambda$ and Gaussian $\lambda$ variants are used to provide dynamic and smoothly varying interpolation weights over multiple preference dimensions, while the unimodal $\lambda$ variant—represented by fixed one-hot vectors such as $[1,0,0,0]$—serves as the baseline where no interpolation is applied, i.e., each training instance aligns with a single fixed objective. This comparison helps assess whether the simplex-weighted combination of preferences brings measurable benefits in robustness, generalization, or alignment quality.
The second set of experiments directly compares $\lambda$-DPO with the original DPO framework. To ensure a fair comparison that isolates the algorithmic difference—rather than the benefit from multiple objectives—this evaluation uses only the unimodal $\lambda$ setting in $\lambda$-DPO, effectively aligning it with single-objective training, just like standard DPO. This design ensures that both models optimize against one fixed preference dimension at a time, allowing us to accurately assess the impact of listwise supervision and policy reparameterization introduced by $\lambda$-DPO.
For the $\lambda$ schedule in total, we therefore investigate three variants:
The random $\lambda$ variant samples a different vector from the simplex for each instance, introducing stochastic diversity across training and encouraging robustness.
The Gaussian $\lambda$ variant samples values from a standard normal distribution and normalizes them to lie on the simplex, enabling smooth but non-uniform emphasis over dimensions.
The unimodal $\lambda$ variant uses fixed one-hot vectors to eliminate interpolation and isolate individual preference components.
Together, these ablation studies provide insights into the roles of interpolation, multi-dimensional preference modeling, and the structural benefits of our listwise optimization strategy.
Conclusion
We presented Multi-Preference Lambda-weighted Listwise DPO, a generalization of Direct Preference Optimization (DPO) that enables alignment of language models with diverse and dynamic human preferences. By extending DPO from a pairwise to a listwise formulation and incorporating multiple preference dimensions via a lambda-weighted aggregation, our approach captures richer supervision signals and allows for controllable alignment without retraining.
Unlike traditional DPO, which assumes a single binary comparison per training instance, Lambda-DPO models the full distribution over ranked outputs, supporting scenarios where user intent is nuanced, multi-objective, or even conflicting. This enables practical applications such as personalized assistants that can adapt to individual user trade-offs (e.g., prioritizing helpfulness over conciseness), regulatory-sensitive deployments that balance safety and informativeness, and modular alignment pipelines where preferences may evolve over time.
Theoretically, our method is grounded in a KL-regularized RL objective, and empirically, we demonstrate competitive or superior performance to standard DPO across alignment benchmarks. By reducing estimation variance and providing interpretable preference control via the lambda vector, Lambda-DPO offers a scalable and expressive solution for real-world preference alignment.
In future work, we plan to explore more structured lambda scheduling, integration with curriculum learning, and extensions to conversational or multimodal alignment tasks.
Acknowledgements
We would like to thank the contributors of the original DPO framework for their foundational work, which inspired this extension. Special thanks to the Hugging Face team and the LLaMA-Factory project for providing robust tools and open-source infrastructure for preference-based fine-tuning. We also acknowledge valuable discussions with colleagues and reviewers who provided feedback during the early stages of this work. Their insights have significantly improved the clarity and rigor of our theoretical analysis.
This work was supported in part by academic research funding and compute resources generously provided by [Your Institution Name Here].
Appendix A A Mathematical Derivations
A.1 Deriving the Final Training Objective
We begin by defining the base reward model used in DPO. Let $\pi_{\theta}(y|x)$ denote the current policy and $\pi_{\text{ref}}(y|x)$ a fixed reference policy. Given input $x$ and candidate outputs $\{y_{1},\ldots,y_{N}\}$, the reward is defined via KL divergence as:
$r_{\theta}(x,y)=\beta\log\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)+\beta\log Z(x)$ (17)
where $Z(x)$ is a normalizing constant ensuring scale-invariance:
$Z(x)=\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)$ (18)
The predicted distribution over completions is then:
$P_{\theta}(y_{i}|x)=\frac{\exp(r_{\theta}(x,y_{i}))}{\sum_{j=1}^{N}\exp(r_{\theta}(x,y_{j}))}$ (19)
Given a human-provided listwise preference distribution $p^{*}_{i}$, we minimize the cross-entropy between $p^{*}_{i}$ and $P_{\theta}$:
$\mathcal{L}^{\text{DPO}}(\theta)=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\exp(r_{\theta}(x,y_{i}))}{\sum_{j=1}^{N}\exp(r_{\theta}(x,y_{j}))}\right)\right]$ (20)
Substituting the reward expression and simplifying:
$\mathcal{L}^{\text{DPO}}(\theta)&=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}Z(x)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}Z(x)^{\beta}}\right)\right]\\ &=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}\right)\right]\end{aligned}$
This becomes the canonical DPO loss. To extend this to multiple preference dimensions, we assume there are $m$ reward components indexed by $k$, each with its own preference distribution $p_{i}^{*(k)}$:
$\mathcal{L}^{(k)}(\theta)=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}\right)\right]$ (21)
We then define a user-controllable or sampled weight vector $\lambda\in\Delta^{m}$ (the $m$-dimensional probability simplex: $\lambda_{k}\geq 0$, $\sum_{k}\lambda_{k}=1$), and define the aggregated loss:
$\mathcal{L}^{\lambda}(\theta)=\sum_{k=1}^{m}\lambda_{k}\cdot\mathcal{L}^{(k)}(\theta)$ (22)
To enable generalization over arbitrary preferences, we optimize the expected loss over $\lambda\sim\text{Unif}(\Delta^{m})$:
$\mathcal{L}_{\text{final}}(\theta)=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\mathcal{L}^{\lambda}(\theta)\right]=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\sum_{k=1}^{m}\lambda_{k}\cdot\mathcal{L}^{(k)}(\theta)\right]$ (23)
A.2 From Final Objective to Lambda-DPO Form
To convert Eq. (25) to a form that operates on distributions directly, we change the order of summation. Recall:
$\mathcal{L}^{(k)}(\theta)=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (24)
So the aggregated expected loss becomes:
$\mathcal{L}_{\text{final}}(\theta)$ $=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\sum_{k=1}^{m}\lambda_{k}\cdot\mathbb{E}_{(x,\{y_{i}\})}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\log P_{\theta}(y_{i}|x)\right]\right]$ (25)
$=\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[-\sum_{i=1}^{N}\sum_{k=1}^{m}\lambda_{k}p_{i}^{*(k)}\log P_{\theta}(y_{i}|x)\right]$ (26)
We now define the aggregated distribution over outputs:
$p^{\lambda}(y_{i}|x)=\sum_{k=1}^{m}\lambda_{k}p^{*(k)}(y_{i}|x)$ (27)
So the objective becomes:
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (28)
Thus, by switching the summation order (i.e., aggregating over preference weights before computing loss), we arrive at an equivalent but more compact objective that is easier to optimize and sample in practice.
A.3 Gradient of Lambda-DPO Loss
We compute the gradient of $\mathcal{L}_{\lambda\text{-DPO}}$:
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}$ $=-\mathbb{E}_{x,\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\nabla_{\theta}\log P_{\theta}(y_{i}\mid x)\right]$ (29)
$=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$ (30)
where we use:
$\nabla_{\theta}P_{\theta}(y_{i}|x)=P_{\theta}(y_{i}|x)\cdot\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)$ (31)
This completes the derivation of the Lambda-weighted DPO objective and its gradient form.
Appendix B B Implementation of _compute_lambda_dpo_loss
This function sequentially performs the following steps:
B.1 Input Tensors and Chunked Log-Probabilities
  1. 1.
    Load the batch tensors:
[…]
    input_ids = inputs["input_ids"]
    attention_mask = inputs["attention_mask"]
    labels = inputs["labels"]
[…]
  2. 2.
    Set the chunk size and compute token/sequence log-probabilities:
    To reduce memory usage, we split the batch into chunks of size $C$, denoted by:
[…]
    chunk_size = self.lambda_dpo_chunk_size or input_ids.size(0)
    seq_log_probs_list = []
    ref_seq_log_probs_list = []
    for start in range(0, input_ids.size(0), chunk_size):
    end = start + chunk_size
    chunk_ids = input_ids[start:end]
    chunk_mask = attention_mask[start:end]
    chunk_labels = labels[start:end]
    logits = model(input_ids=chunk_ids, attention_mask=chunk_mask).logits
    log_probs = F.log_softmax(logits[:, :-1], dim=-1)
    labels_shifted = chunk_labels[:, 1:].clone()
    mask = labels_shifted != self.label_pad_token_id
    labels_shifted[labels_shifted == self.label_pad_token_id] = 0
    label_log_probs = torch.gather(log_probs, dim=2, index=labels_shifted.unsqueeze(-1)).squeeze(-1)
    seq_lp = (label_log_probs * mask).sum(dim=1) / mask.sum(dim=1)
    seq_log_probs_list.append(seq_lp)
  3. 3.
    Compute reference log-probabilities (if applicable):
    $\log\pi_{\text{ref}}(y|x)=\begin{cases}\log\text{softmax}(\text{ref\_model}(x)),&\text{if ref model is used}\\ 0,&\text{otherwise}\end{cases}$
    with torch.no_grad():
    if self.finetuning_args.use_ref_model and self.ref_model is not None:
    ref_logits = self.ref_model(input_ids=chunk_ids, attention_mask=chunk_mask).logits
    ref_log_probs = F.log_softmax(ref_logits[:, :-1], dim=-1)
    ref_label_log_probs = torch.gather(ref_log_probs, dim=2, index=labels_shifted.unsqueeze(-1)).squeeze(-1)
    ref_lp = (ref_label_log_probs * mask).sum(dim=1) / mask.sum(dim=1)
    else:
    ref_lp = torch.zeros_like(seq_lp)
    ref_seq_log_probs_list.append(ref_lp)
    seq_log_probs = torch.cat(seq_log_probs_list, dim=0)
    ref_seq_log_probs = torch.cat(ref_seq_log_probs_list, dim=0)
B.2 Reference Model and Relative Reward
  1. 3.
    Compute the relative reward:
    $r_{\theta}(x,y)=\beta\left(\log\pi_{\theta}(y|x)-\log\pi_{\text{ref}}(y|x)\right)$
    Then reshape:
    $r_{\theta}\in\mathbb{R}^{B\times 4\times 4},\quad\pi^{*}\in\mathbb{R}^{B\times 4\times 4}$
    r_theta = self.beta * (seq_log_probs - ref_seq_log_probs)
    B = input_ids.size(0) // 16
    r_theta = r_theta.view(B, 4, 4)
    pi_target = pi_target.view(B, 4, 4)
B.3 Listwise Loss by Preference Dimension
  1. 5.
    For each preference axis $i\in\{1,2,3,4\}$, compute the listwise loss:
    $\ell^{(i)}=-\sum_{j}p^{*(i)}_{j}\cdot\log\left(\frac{\exp(r^{(i)}_{j})}{\sum_{j^{\prime}}\exp(r^{(i)}_{j^{\prime}})}\right)$
    listwise_losses = []
    for i in range(4):
    r = r_theta[:, i, :]
    p = pi_target[:, i, :] + 1e-10
    log_softmaxed = r - torch.logsumexp(r, dim=1, keepdim=True)
    loss = -(p * log_softmaxed).sum(dim=1).mean()
    listwise_losses.append(loss)
B.4 Lambda Weights and Final Loss
  1. 6.
    Sample a random weight vector $\boldsymbol{\lambda}\sim\Delta^{4}$, and compute:
    $\mathcal{L}(\theta)=\sum_{i=1}^{4}\lambda_{i}\ell^{(i)}$
    weights = torch.rand(4, device=input_ids.device)
    weights /= weights.sum()
    final_loss = sum(w * l for w, l in zip(weights, listwise_losses))
    if return_outputs:
    return final_loss, {"loss_components": listwise_losses}
    return final_loss
Appendix C: Humorous Overview of the Added Classes
C.1 UltrafeedbackDatasetConverter
Input
A single UltraFeedback record. It contains:
  • a user prompt (instruction)
  • multiple completions, each with four ratings: helpfulness, honesty, instruction_following, and truthfulness.
Output
A dictionary with the prompt and, for every dimension, the raw scores plus a preference vector.
Example
Input:
instruction = "Explain photosynthesis."
completions = [
{"response": "...", "annotations": {"helpfulness": {"Rating": 4}}},
{"response": "...", "annotations": {"helpfulness": {"Rating": 2}}},
{"response": "...", "annotations": {"helpfulness": {"Rating": 5}}},
{"response": "...", "annotations": {"helpfulness": {"Rating": 3}}}
Output (simplified):
"_prompt": [{"role": "user", "content": "Explain photosynthesis."}],
"helpfulness": [
{"model": "?", "content": "...", "score": 4},
"_pi_target": {"helpfulness": [0.24, 0.09, 0.45, 0.22]}
C.2 ListwiseDatasetProcessor
Input
A dictionary produced by the converter that contains _prompt, the lists of responses for each dimension, and _pi_target.
Output
A flattened structure ready for training: for every sample there are 16 sequences of token IDs (one per response per dimension), matching labels and attention masks, and the list of probabilities.
Example
Input:
"_prompt": ...,
"helpfulness": [
{"content": "A"}, {"content": "B"}, {"content": "C"}, {"content": "D"}
"_pi_target": {"helpfulness": [0.24, 0.09, 0.45, 0.22], ...}
Output (illustrative shapes):
"input_ids": [seq1, seq2, ..., seq16], # each seq corresponds to a prompt+response
"labels": [...], # same layout as input_ids
"attention_mask": [...], # masks per sequence
"pi_target": [0.24, 0.09, 0.45, 0.22, ...] # flattened list of probabilities
C.3 ListwiseDataCollatorWithPadding
Input
A list of processed samples where each sample holds 16 sequences of token IDs and the flattened probability list.
Output
A padded batch containing tensors input_ids, attention_mask, labels, and an accompanying tensor of preference values aligned with the responses.
Example
Input:
# List of samples where each sample holds 16 sequences and
# a corresponding 16-element probability list
[example1, example2, ..., example_B] # B = batch size
Output:
"input_ids": Tensor of shape [B * 16, L],
"attention_mask": Tensor of shape [B * 16, L],
"labels": Tensor of shape [B * 16, L],
"pi_target": Tensor of shape [B * 16]
These three classes transform UltraFeedback data into tensors suitable for Lambda-weighted DPO training, while preserving the four-response structure and converting ratings into softmax preference vectors.
This function performs sequence probability calculation, reference model comparison, listwise cross-entropy over preference dimensions, and $\lambda$-weighted aggregation. It fully realizes the lambda-DPO loss objective described in the paper’s methodology section.