### v4

# Multi-Preference Lambda-weighted Listwise DPO for Dynamic Preference Alignment

Abstract
While large language models (LLMs) excel at text generation, aligning them with human preferences remains challenging. Reinforcement learning from human feedback (RLHF) improves alignment but is costly and unstable. Direct Preference Optimization (DPO) offers a simpler alternative, yet assumes a fixed, single-dimensional preference.
We propose Multi-Preference Lambda-weighted Listwise DPO, a generalization of DPO that supports multiple preference dimensions and dynamic interpolation via a simplex-weighted $\lambda$ vector. Our method enables listwise supervision and flexible alignment without re-training. While our experiments are conducted on 1B–2B scale models, this is an intentional choice: smaller models provide a more stringent testbed where performance improvements more clearly reflect the effectiveness of the alignment strategy itself. Moreover, such models are widely used in compute-constrained applications, making our improvements both methodologically meaningful and practically valuable. Empirical results show that our approach matches or surpasses standard DPO on alignment benchmarks while offering improved adaptability.
Introduction
Large language models (LLMs) have achieved impressive results across natural language tasks, but often misalign with human preferences—producing responses that are factually incorrect, verbose, or unhelpful. This misalignment stems from the next-token prediction objective, which lacks explicit human feedback.
To address this, recent approaches incorporate preference supervision. RLHF trains a reward model from human comparisons and optimizes it via reinforcement learning, but suffers from high cost, instability, and deployment complexity. Direct Preference Optimization (DPO) offers a simpler alternative by directly optimizing model outputs to reflect pairwise preferences, achieving strong results with lower training overhead. However, DPO assumes a single reward dimension and cannot adapt to varying or multi-faceted user intents.
In practice, alignment involves multiple criteria—e.g., helpfulness, honesty, fluency—whose importance may shift across tasks. To address this, we propose Multi-Preference Lambda-weighted Listwise DPO, a generalization of DPO that incorporates multiple preference axes and allows dynamic control via a weight vector $\lambda$ drawn from the simplex.
Our method extends DPO from pairwise to listwise supervision and introduces a learnable scheduling mechanism over $\lambda$ vectors, enabling flexible, fine-grained alignment. Empirical results on multi-objective benchmarks show that our method matches or exceeds standard DPO while offering greater control and adaptability.
Related Work
RLHF and DPO.
Reinforcement learning from human feedback (RLHF) [18, 1] aligns LLMs via reward models and PPO [21], but is costly and unstable [28, 5]. Direct Preference Optimization (DPO) [19] simplifies alignment by reframing it as binary classification on preference pairs, improving stability and efficiency. Extensions like Quark [26] and LLaMA-DPO [16] show strong performance, though limited to fixed, pairwise objectives.
Listwise Preference Learning.
Listwise learning captures global ranking better than pairwise [4, 24]. UltraFeedback [8] provides ranked human feedback, and methods like GLOW [25] train reward models from such data. Our approach differs by optimizing listwise loss directly in policy space, without intermediate reward modeling.
Multi-Objective and Controllable Alignment.
Prior work on multi-reward RL [12], decomposition [13], and reward interpolation [23] requires retraining for each preference trade-off. Our method incorporates a weight vector $\lambda$ during training, enabling generalization to arbitrary preferences. Related ideas include universal value function approximators [20]. Steerable LMs using prompts [3, 22], control tokens [11], or constrained decoding [17] offer limited flexibility. In contrast, our method enables fine-grained, multi-dimensional alignment via $\lambda$ without modifying prompts or decoding.
Preliminaries
We begin by reviewing the foundation of Direct Preference Optimization (DPO) and its connection to KL-regularized reinforcement learning from human feedback (RLHF). Our setting assumes the existence of a base language model $\pi_{\text{ref}}$, a supervised fine-tuned (SFT) model, and a collection of ranked or pairwise human preferences. We aim to train a target policy $\pi_{\theta}$ that better aligns with human preferences while remaining close to $\pi_{\text{ref}}$.
Pairwise Preference Modeling.
Let $x$ denote the prompt, and $y_{w}\succ y_{l}$ represent a human preference between a pair of completions: $y_{w}$ (the preferred answer) and $y_{l}$ (the less preferred one). These preferences are sampled from an underlying distribution $p^{*}(y_{w},y_{l}|x)$, which is typically assumed to be induced by a latent reward function $r(x,y)$. A commonly adopted model is the Bradley-Terry (BT) model [2]:
$p^{*}(y_{w}\succ y_{l}\mid x)=\frac{\exp(r(x,y_{w}))}{\exp(r(x,y_{w}))+\exp(r(x,y_{l}))}$ (1)
Given a dataset of comparisons $\mathcal{D}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$, we can learn the reward function $r(x,y)$ by minimizing the negative log-likelihood under the BT model:
$\mathcal{L}_{\text{BT}}(r)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma(r(x,y_{w})-r(x,y_{l}))\right],$ (2)
where $\sigma(z)=1/(1+\exp(-z))$ is the logistic sigmoid.
KL-Regularized RL Objective.
Once a reward model is trained, the standard RLHF approach fine-tunes the policy $\pi_{\theta}$ to maximize expected rewards while constraining deviation from the reference policy $\pi_{\text{ref}}$ via KL regularization:
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}\left[r(x,y)-\beta\cdot\mathrm{KL}\left(\pi_{\theta}(y|x)\|\pi_{\text{ref}}(y|x)\right)\right],$ (3)
where $\beta>0$ controls the strength of the KL constraint.
This objective encourages the policy to produce high-reward outputs while staying close to $\pi_{\text{ref}}$. However, in practice, it requires sampling, reward evaluation, and reinforcement learning algorithms such as PPO [21] to optimize.
Reparameterization and DPO.
DPO sidesteps explicit reward modeling and reinforcement learning by recognizing that the optimal policy $\pi_{r}$ under Equation 3 can be written as:
$\pi_{r}(y|x)\propto\pi_{\text{ref}}(y|x)\cdot\exp\left(\frac{1}{\beta}r(x,y)\right).$ (4)
Using Equation 4, we reparameterize the reward difference into the log-ratio of policies:
$\beta(r(x,y_{w})-r(x,y_{l}))=\log\frac{\pi_{r}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{r}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}.$ (5)
Replacing $\pi_{r}$ with $\pi_{\theta}$ yields the DPO loss:
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right].$ (6)
This loss directly optimizes the policy parameters $\theta$ using preference comparisons, without sampling or reinforcement learning. It forms the foundation upon which we build our multi-preference and listwise extensions in later sections.
Multi-Preference Lambda-weighted Listwise DPO
Motivated by the limitations of pairwise DPO under a single fixed preference distribution, we introduce a generalized framework that supports listwise preference supervision across multiple human objectives. In real-world settings, human judgments often reflect diverse criteria—such as helpfulness, harmlessness, and conciseness—and capturing these via a unified formulation enables controllable alignment across axes of interest.
Our key insight is that we can reparameterize multi-dimensional listwise preferences using a convex combination of multiple reward signals. Specifically, we define a simplex-weighted listwise loss that extends the DPO framework to dynamic, multi-objective optimization.
Listwise Preference Modeling.
Let $\{y_{1},y_{2},\dots,y_{N}\}$ be a set of candidate completions for a prompt $x$, and let $p^{*}(y_{i}|x)$ be a listwise human preference distribution over these candidates. Analogous to the pairwise DPO reparameterization, we define the listwise policy distribution under $\pi_{\theta}$ as:
$P_{\theta}(y_{i}|x)=\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}},$ (7)
which defines a softmax over reweighted log-ratios, consistent with KL-derived optimal policies.
We then define the listwise DPO objective as a cross-entropy loss between human preferences $p^{*}(y_{i}|x)$ and model predictions:
$\mathcal{L}_{\text{listwise}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[\sum_{i=1}^{N}p^{*}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right].$ (8)
Lambda-weighted Preference Composition.
To enable multi-objective control, we assume access to $m$ distinct preference dimensions, each defined by a distribution $p^{*(k)}(y_{i}|x)$ for objective $k\in\{1,\dots,m\}$. For each prompt, we sample or specify a weight vector $\lambda\in\Delta^{m}$ (the probability simplex), and define the aggregated preference distribution:
$p^{\lambda}(y_{i}|x)=\sum_{k=1}^{m}\lambda_{k}\cdot p^{*(k)}(y_{i}|x).$ (9)
This allows us to interpolate among various alignment goals without retraining.
The full Lambda-weighted Listwise DPO objective becomes:
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right].$ (10)
This expectation can be approximated by sampling $\lambda\sim\text{Unif}(\Delta^{m})$ or conditioning on user-specified preferences.
What does $\lambda$-DPO update do?
To understand the optimization dynamics, consider the gradient of the loss:
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}|x)-P_{\theta}(y_{i}|x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\right].$ (11)
This gradient updates the model to increase likelihood of responses aligned with weighted human preferences, while decreasing likelihood of responses overrepresented in the reference policy. Importantly, when $\lambda$ is fixed (e.g., $\lambda=(1,0,0)$), the objective reduces to a single-objective listwise DPO loss.
Lambda-DPO Pipeline.
The training process proceeds as follows:
  1. 1.
    For each prompt $x$, obtain $N$ candidate completions $\{y_{1},\dots,y_{N}\}\sim\pi_{\text{ref}}(\cdot|x)$.
  2. 2.
    For each alignment dimension $k$, elicit a human listwise preference distribution $p^{*(k)}(y_{i}|x)$.
  3. 3.
    Sample or define $\lambda\in\Delta^{m}$, and compute the mixture preference $p^{\lambda}$.
  4. 4.
    Optimize $\pi_{\theta}$ using the listwise cross-entropy loss between $p^{\lambda}$ and $P_{\theta}$.
In practice, we can reuse datasets like UltraFeedback [8], where multiple preference criteria are annotated. During inference, changing $\lambda$ enables personalized alignment without modifying model parameters.
Theoretical Analysis of Multi-Preference Lambda-weighted Listwise DPO
While Direct Preference Optimization (DPO) offers a robust framework for aligning language models with human preferences via pairwise supervision, it inherently simplifies the preference space by assuming binary feedback between two completions. In this section, we analyze the theoretical implications of this simplification and motivate the use of a listwise and multi-preference extension.
Pairwise DPO and Its Limitations.
Given a prompt $x$ and two outputs $y_{w}$ (winner) and $y_{l}$ (loser), DPO defines a binary preference distribution:
$P^{*}(y_{i}|x)=\begin{cases}1&\text{if }y_{i}=y_{w}\\ 0&\text{if }y_{i}=y_{l}\end{cases}$ (12)
This implicitly constrains the learning objective to satisfy $P^{*}(y_{w}|x)+P^{*}(y_{l}|x)=1$, ignoring the quality of other plausible candidates. Consequently, the optimized policy $\pi_{\theta}$ may fail to capture the full spectrum of user preferences, introducing estimation error.
Listwise Distribution via Softmax.
To mitigate this, we define a listwise target distribution $\pi^{*}$ over all completions $\{y_{1},y_{2},\dots,y_{N}\}$ for prompt $x$, derived from scalar reward scores $r(x,y_{i})$:
$p^{*}(y_{i}\mid x)=\frac{\exp(r(x,y_{i}))}{\sum_{j=1}^{N}\exp(r(x,y_{j}))}$ (13)
This distribution captures relative preference over all candidates, allowing for more representative updates and lower variance.
Multi-Preference Supervision.
Real-world alignment involves multiple competing objectives, such as helpfulness, safety, and factuality. Let $p^{*(k)}(y_{i}|x)$ denote the listwise preference under objective $k\in\{1,\dots,m\}$. We define a convex combination of these objectives via a preference vector $\lambda\in\Delta^{m}$:
$p^{\lambda}(y_{i}\mid x)=\sum_{k=1}^{m}\lambda_{k}\cdot p^{*(k)}(y_{i}\mid x)$ (14)
The training objective minimizes the cross-entropy between the aggregated distribution $p^{\lambda}$ and the policy-induced distribution $P_{\theta}$:
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (15)
Gradient Interpretation.
The gradient of this objective is:
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$ (16)
This update rule drives $\pi_{\theta}$ toward the weighted human preference distribution $p^{\lambda}$, allowing alignment to reflect fine-grained and multi-dimensional values.
Variance Reduction via Listwise Supervision.
Using a normalized softmax distribution $\pi^{*}$ derived from reward scores avoids the instability and high variance that arise when optimizing based on hard binary labels. Empirically, listwise training improves sample efficiency and convergence by utilizing the full output ranking structure.
Summary.
In summary, Multi-Preference Lambda-weighted Listwise DPO:
  • Recovers a more informative supervision signal via $\pi^{*}$.
  • Reduces estimation error and variance in gradient updates.
  • Enables alignment over a continuous space of user-defined objectives via $\lambda$.
Experiment Setup
Experimental Objectives and Comparison Targets
The primary goal of our experiments is to evaluate the performance of the proposed $\lambda$-DPO algorithm compared to the standard DPO method. Specifically, we assess their effectiveness across different model architectures and training settings using public preference datasets and standard evaluation benchmarks.
We consider two levels of comparison:
Model comparison: We experiment with two strong open-source language models—LLaMA and Qwen. These models differ in architecture, pretraining corpora, and tokenizer, providing a robust testbed for evaluating generalizability.
Method comparison: We compare DPO and $\lambda$-DPO. For each method and model pair, we perform independent hyperparameter tuning (e.g., learning rate, batch size, $\beta$) to ensure that comparisons are fair and reflect the best achievable performance under comparable resource constraints.
Training Settings
Inspired by the SimPO baseline setup, we adopt two training pipelines for both models:
  • Base: The model is directly trained on the UltraFeedback dataset using DPO or $\lambda$-DPO. No supervised fine-tuning is applied before preference optimization. This setting isolates the influence of the preference learning method alone.
  • Finetuned-Base: The model is first supervised fine-tuned (SFT) on the UltraChat-200k dataset, which contains approximately 200K high-quality dialogue examples. The resulting SFT model is then optimized using UltraFeedback preferences via DPO or $\lambda$-DPO. This pipeline reflects realistic alignment procedures.
All training is conducted on A100 GPUs, with consistent batch sizes and sequence lengths. For each setting, hyperparameters are tuned via validation performance on reward preference accuracy and KL stability.
Evaluation Benchmarks and Metrics
We evaluate model performance using the lm-eval-harness [9] framework, which tests accuracy on a diverse set of benchmarks covering reasoning, factual correctness, and commonsense understanding.
Specifically, we report average accuracy across the following six established datasets:
  • MMLU [10]: 57 subject areas with around 15K test questions targeting academic and professional knowledge.
  • ARC (Challenge) [6]: Grade-school science questions ($\sim$1.1K) requiring multi-step reasoning.
  • HellaSwag [27]: 10K commonsense sentence completion tasks with adversarial distractors.
  • TruthfulQA [15]: 817 questions designed to expose hallucinations and factual errors.
  • Winograd [14]: 273 examples testing pronoun coreference and syntactic disambiguation.
  • GSM8K [7]: 8.5K arithmetic reasoning problems with step-by-step solutions.
Ablation Experiment
To better understand the impact of the $\lambda$ vector in $\lambda$-DPO, we conduct ablation experiments focusing on how different preference weightings affect alignment performance.
Specifically, we compare two $\lambda$ schedule variants within the listwise $\lambda$-DPO framework:
[…]
  • Unimodal $\lambda$: we use fixed one-hot vectors (e.g., $(1,0,0,\dots)$), effectively disabling interpolation and isolating optimization to a single preference dimension per instance. This variant serves as a control version without multi-dimensional aggregation.
This comparison helps isolate the contribution of $\lambda$-based preference interpolation within the listwise setting. For a separate comparison between standard DPO and our listwise formulation, please refer to the main benchmark results in Table 1.
Comparative Experiment
We compare standard DPO and our proposed $\lambda$-DPO under identical training and evaluation setups on two models: LLaMA 3.2B Instruct and Qwen 1.8B Chat. Both models are first supervised fine-tuned (SFT) on UltraChat-200k, then aligned using the UltraFeedback listwise dataset.
Training Details.
We use LoRA (rank 8, all-layer target) with a learning rate of $5\times 10^{-6}$, batch size 1, gradient accumulation steps 8, cosine learning rate schedule, and 5 epochs. The KL penalty $\beta$ is set to 0.1. $\lambda$-DPO uses the lambda_dpo loss with chunk size 1.
Evaluation.
We evaluate on six benchmarks (MMLU, ARC, HellaSwag, TruthfulQA, WSC273, GSM8K) using the lm-eval-harness, reporting accuracy as the primary metric.
Main Results
Table 1 reports the accuracy of DPO and $\lambda$-DPO on six representative benchmarks for both LLaMA 3.2B Instruct and Qwen 1.8B Chat. On LLaMA, $\lambda$-DPO consistently improves performance across all tasks, with notable gains such as +1.1% on ARC, +3.3% on GSM8K, and +3.6% on TruthfulQA. The improvement on WSC273 is also evident (+1.8%). For Qwen 1.8B Chat, $\lambda$-DPO outperforms DPO on all benchmarks except TruthfulQA, where DPO achieves a slightly higher score. Notably, large gains are observed on GSM8K (+14.7%) and MMLU (+3.9%), demonstrating the generalization benefits of $\lambda$-DPO across different model families.
Model Method MMLU ARC HellaSwag TruthfulQA WSC273 GSM8K
LLaMA 3.2B Instruct DPO 0.458 0.364 0.456 0.378 0.696 0.312
$\lambda$-DPO 0.460 0.375 0.459 0.414 0.714 0.345
Qwen 1.8B Chat DPO 0.379 0.271 0.418 0.366 0.656 0.008
$\lambda$-DPO 0.418 0.284 0.428 0.312 0.667 0.155
Table 1: Accuracy comparison between DPO and $\lambda$-DPO on six benchmarks using LLaMA 3.2B Instruct and Qwen 1.8B Chat. Bold values indicate better performance for each benchmark.
Visualization
To further illustrate these results, we provide bar charts comparing DPO and $\lambda$-DPO performance on the two models.
Figure 1: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B Instruct.
Figure 2: Accuracy comparison between DPO and $\lambda$-DPO on Qwen 1.8B Chat.
Performance Analysis and Discussion
As shown in both Table 1 and Figures 12, $\lambda$-DPO consistently improves performance across most tasks. The method yields notable improvements on ARC, GSM8K, and MMLU, while maintaining parity on Winogrande. TruthfulQA is the only benchmark where DPO slightly outperforms $\lambda$-DPO (in the Qwen model).
These results suggest that $\lambda$-DPO effectively captures human preference signals across diverse task types. Its listwise preference modeling contributes to better alignment with human-evaluated quality metrics such as factuality, helpfulness, and fluency.
Nonetheless, for certain adversarial or logic-heavy tasks like TruthfulQA, the benefits of listwise supervision may be less reliable. Integrating $\lambda$-DPO with structured supervision—such as chain-of-thought prompting [22]—could potentially mitigate this gap.
Conclusion of Comparative Analysis
In summary, $\lambda$-DPO demonstrates consistent improvements over DPO in most evaluation scenarios, confirming the effectiveness of listwise preference supervision. Its generalization advantage makes it a strong candidate for broader alignment tasks. Future work may explore hybrid strategies that combine $\lambda$-DPO with structured reasoning methods to further enhance performance on complex benchmarks.
Ablation Experiment
To further examine the effectiveness of different preference vector settings, we conduct an ablation experiment focusing on four unimodal $\lambda$-DPO models. Each model is trained using a fixed one-hot $\lambda$ vector, targeting only one of the four annotated dimensions: helpfulness, honesty, instruction-following, and truthfulness. This setup allows us to isolate the influence of each individual preference axis and compare their respective impacts across evaluation benchmarks.
Unimodal Comparison Across Benchmarks
We evaluate the four unimodal $\lambda$-DPO variants across six standard benchmarks—mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k. Figure 9 shows the accuracy distribution of each model across these tasks.
Figure 3: Accuracy Comparison of Unimodal $\lambda$-DPO Models with Fixed Preference Vectors.
We observe that all four models achieve relatively comparable performance across the six benchmarks. The $\lambda=[0,1,0,0]$ model (favoring honesty) achieves the highest average accuracy, while the $\lambda=[1,0,0,0]$ model (favoring helpfulness) performs best on truthfulqa and shows strong performance on gsm8k. The models focusing on instruction-following and fluency show only marginally lower scores, with no major degradation in benchmark performance.
Preference Ranking and Vector Hypothesis
To better assess how each individual preference dimension contributes to generalization performance, we compute the average accuracy of the four unimodal $\lambda$-DPO models across six benchmark tasks:
  • $\lambda=[1,0,0,0]$: 0.4593
  • $\lambda=[0,1,0,0]$: 0.4598
  • $\lambda=[0,0,1,0]$: 0.4543
  • $\lambda=[0,0,0,1]$: 0.4548
  • $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$: 0.4612
The results indicate that all four unimodal preference dimensions contribute to model alignment, with helpfulness and honesty yielding slightly better generalization. However, the uniform interpolation $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ achieves the highest overall accuracy (0.4612), outperforming all unimodal variants.
This highlights the benefit of explicitly retaining the $\lambda$ vector for multi-objective supervision. Unlike standard DPO, which implicitly collapses preferences into a fixed form, $\lambda$-DPO provides fine-grained control over alignment trade-offs and leads to more robust generalization.
Discussion and Ablation Analysis
To further understand how individual preference dimensions influence alignment, we observe from Figure 9 that:
  • The $\lambda=[1,0,0,0]$ (helpfulness) model performs best on truthfulqa and achieves solid results on gsm8k.
  • The $\lambda=[0,1,0,0]$ (honesty) model leads slightly in overall average accuracy, especially strong on mmlu.
  • The $\lambda=[0,0,1,0]$ (instruction-following) and $\lambda=[0,0,0,1]$ (fluency) models show close and competitive performance, contrary to the notion that they are stylistically weaker.
These results indicate that optimizing for either content-driven or stylistic objectives in isolation yields comparable outcomes under our listwise framework. It further motivates the integration of multiple preference signals, especially for tasks requiring both correctness and user-aligned expressiveness.
In future work, we plan to extend this ablation framework by exploring:
  • Adaptive $\lambda$ scheduling, where preference weights change based on input context or task type.
  • Hierarchical modeling that separates semantic utility from stylistic fluency.
  • Persona- or task-specific interpolation strategies to match different downstream alignment needs.
Performance-Driven Lambda Scheduling via Polynomial Modeling
We introduce a general and flexible method for performance-driven lambda scheduling, which constructs a probability distribution over preference vectors $\lambda\in\Delta^{d}$ based on observed alignment performance. Each supervised training point $(\lambda^{(i)},y^{(i)})$ consists of a preference configuration $\lambda^{(i)}$ and a corresponding scalar performance metric $y^{(i)}$ (e.g., average benchmark accuracy). The scheduling process unfolds in four steps:
Step 1: Polynomial Feature Mapping
We begin by mapping each preference vector $\lambda$ to a high-dimensional feature space using a polynomial embedding. Specifically, we define a degree-$p$ polynomial feature map $\phi_{p}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N}$ that enumerates all monomials up to degree $p$:
$\phi_{p}(\lambda)=\left[1,\lambda_{1},\dots,\lambda_{d},\lambda_{1}^{2},\lambda_{1}\lambda_{2},\dots,\lambda_{d}^{p}\right]^{\top},$ (17)
where the total number of features is given by:
$N=\binom{d+p}{p}.$ (18)
This expansion enables expressive yet tractable modeling of non-linear relationships between $\lambda$ vectors and alignment performance.
Step 2: Polynomial Regression
Using the training set $\{(\lambda^{(i)},y^{(i)})\}_{i=1}^{n}$, we fit a least-squares regression model in the lifted feature space to learn the performance landscape:
$\min_{\mathbf{w}\in\mathbb{R}^{N}}\sum_{i=1}^{n}\left(\mathbf{w}^{\top}\phi_{p}(\lambda^{(i)})-y^{(i)}\right)^{2}.$ (19)
The resulting predictive function
$f(\lambda)=\mathbf{w}^{\top}\phi_{p}(\lambda)$ (20)
approximates the expected performance of any candidate $\lambda$ on downstream alignment tasks.
Step 3: Grid Construction over the Simplex
To enable tractable sampling from the $\lambda$ space, we construct a discrete candidate set $\{\lambda^{(j)}\}_{j=1}^{k}\subset\Delta^{d}$ via one of the following methods:
  • Simplex discretization: generate all combinations with replacement and normalize;
  • Dirichlet sampling: draw $\lambda^{(j)}\sim\mathrm{Dir}(\boldsymbol{\alpha})$, where $\boldsymbol{\alpha}=\mathbf{1}$ yields a uniform prior over the simplex.
Step 4: Softmax-Based Scheduling Distribution
We convert the predicted scores $s_{j}=f(\lambda^{(j)})$ into a probability distribution over the candidate $\lambda$ vectors using a softmax transformation:
$p(\lambda^{(j)})=\frac{\exp(\tau\cdot f(\lambda^{(j)}))}{\sum_{l=1}^{k}\exp(\tau\cdot f(\lambda^{(l)}))},$ (21)
where the temperature parameter $\tau>0$ controls the exploration–exploitation trade-off. Higher values of $f(\lambda^{(j)})$ correspond to better expected performance, allowing the scheduler to prioritize promising $\lambda$ vectors while retaining stochasticity. This yields a flexible, differentiable mechanism for guiding multi-objective alignment through learned preference scheduling.
Step 5: Empirical Instantiation with Five Training Points
To validate the polynomial scheduling framework, we use five observed training points—four unimodal $\lambda$ vectors and one uniform configuration—each paired with an empirical alignment accuracy:
$\lambda^{(1)}$ $=[1,0,0,0],\quad y^{(1)}=0.4593$
$\lambda^{(2)}$ $=[0,1,0,0],\quad y^{(2)}=0.4598$
$\lambda^{(3)}$ $=[0,0,1,0],\quad y^{(3)}=0.4543$
$\lambda^{(4)}$ $=[0,0,0,1],\quad y^{(4)}=0.4548$
$\lambda^{(5)}$ $=[0.25,0.25,0.25,0.25],\quad y^{(5)}=0.4612$
We set the polynomial degree to $p=2$, yielding a total of $N=\binom{4+2}{2}=15$ monomial features. Using gradient descent to minimize the squared error over the training set, we obtain the fitted predictive model:
$f(\lambda)$ $=0.35+0.09\lambda_{1}+0.09\lambda_{2}+0.09\lambda_{3}+0.09\lambda_{4}$
$+0.02\lambda_{1}^{2}+0.04\lambda_{1}\lambda_{2}+0.03\lambda_{1}\lambda_{3}+0.02\lambda_{1}\lambda_{4}$
$+0.02\lambda_{2}^{2}+0.02\lambda_{2}\lambda_{3}+0.02\lambda_{2}\lambda_{4}$
$+0.01\lambda_{3}^{2}+0.00\lambda_{3}\lambda_{4}+0.01\lambda_{4}^{2}$
Step 6: Lambda Sampling from Learned Scheduler
We discretize the simplex $\Delta^{4}$ by uniformly sampling $k=10$ candidate $\lambda$ vectors from a Dirichlet$(\boldsymbol{1})$ prior to ensure diverse coverage of the preference space. For each sample $\lambda^{(j)}$, we evaluate its predicted alignment score $s_{j}=f(\lambda^{(j)})$ using the learned polynomial model, and convert these scores into a sampling distribution using softmax:
$p(\lambda^{(j)})=\frac{\exp(\tau\cdot f(\lambda^{(j)}))}{\sum_{l=1}^{10}\exp(\tau\cdot f(\lambda^{(l)}))},\quad\text{with }\tau=100.$
This yields a discrete distribution over candidate $\lambda$ vectors that prioritizes those with higher predicted alignment quality. The table below summarizes the 10 sampled vectors, their predicted scores, and their corresponding softmax probabilities:
$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $f(\lambda)$ $p(\lambda)$
0.212 0.334 0.245 0.209 0.452 0.095
0.126 0.237 0.131 0.507 0.450 0.075
0.542 0.079 0.256 0.123 0.455 0.117
0.233 0.721 0.020 0.025 0.459 0.185
0.004 0.334 0.281 0.381 0.449 0.067
0.320 0.069 0.513 0.099 0.451 0.082
0.236 0.155 0.358 0.251 0.452 0.093
0.141 0.110 0.701 0.048 0.454 0.110
0.070 0.198 0.388 0.344 0.451 0.081
0.090 0.139 0.306 0.465 0.450 0.095
Table 2: Sampled $\lambda$ vectors from Dirichlet(1,1,1,1)$1$(1,1,1,1)( 1 , 1 , 1 , 1 ), their predicted scores $f(\lambda)$, and softmax-based sampling probabilities $p(\lambda)$ with temperature $\tau=100$.
Step 7: Scheduler vs. Uniform Comparison
To assess the overall utility of the learned scheduling distribution, we compare it against a baseline that uses a fixed uniform preference vector $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ for all training examples. Both approaches are applied under identical listwise $\lambda$-DPO training conditions using the LLaMA 3.2B Instruct model.
Figure 4 shows the empirical accuracy across six alignment benchmarks. The learned scheduler achieves a slightly higher average accuracy (0.462 vs. 0.461), validating its effectiveness as a data-driven alternative to static preference weighting.
Beyond average performance, the learned scheduler offers several additional advantages: (i) it introduces sample-level variability by stochastically selecting from a high-quality preference distribution, which can act as an implicit regularizer during training; and (ii) it enables better coverage of the preference simplex, reducing the risk of overfitting to limited preference configurations and helping mitigate modeling bias in the polynomial predictor.
Figure 4: Comparison between learned $\lambda$ scheduler and uniform $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ across six alignment benchmarks. The learned scheduler achieves a slightly higher average accuracy and improves robustness by distributing preference sampling over higher-quality regions of the simplex.
Conclusion
We introduced Multi-Preference Lambda-weighted Listwise DPO, which extends DPO to a listwise setting with controllable $\lambda$ vectors for multi-dimensional alignment. This approach enables flexible, interpretable optimization over preference trade-offs without retraining.
Experiments on 1B–2B scale models show consistent gains over standard DPO, with average accuracy improvements of +1.7% on LLaMA 3.2-1B Instruct and +2.8% on Qwen-1.8 Chat. We further propose a learned scheduler that adaptively samples high-performing $\lambda$ configurations, outperforming uniform baselines while reducing modeling bias. These results highlight the effectiveness of listwise, preference-aware alignment in instruction tuning—especially for efficiently improving smaller models where preference supervision has greater impact, capacity is limited, and retraining is impractical. This is particularly relevant for domain-specific assistants on edge devices, multilingual student models distilled from larger systems, or rapid adaptation via methods like LoRA. In such settings, dynamically adjusting alignment objectives—e.g., helpfulness vs. honesty—without retraining offers both efficiency and controllability.
Acknowledgements
We thank the contributors of the original DPO framework for their foundational work, which inspired this extension. We also appreciate the open-source efforts of the Hugging Face team and the LLaMA-Factory project for enabling scalable preference-based fine-tuning.
This work was made possible through collaborative effort and self-supported compute resources by the authors: Yuhui Sun, Xiyao Wang, Zixi Li, Zhenlong Yuan and Jinman Zhao. We further thank colleagues and early readers for helpful feedback that improved the theoretical formulation and presentation.
Appendix A A Mathematical Derivations
A.1 Deriving the Final Training Objective
We begin by defining the base reward model used in DPO. Let $\pi_{\theta}(y|x)$ denote the current policy and $\pi_{\text{ref}}(y|x)$ a fixed reference policy. Given input $x$ and candidate outputs $\{y_{1},\ldots,y_{N}\}$, the reward is defined via KL divergence as:
$r_{\theta}(x,y)=\beta\log\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)+\beta\log Z(x)$ (22)
where $Z(x)$ is a normalizing constant ensuring scale-invariance:
$Z(x)=\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)$ (23)
The predicted distribution over completions is then:
$P_{\theta}(y_{i}|x)=\frac{\exp(r_{\theta}(x,y_{i}))}{\sum_{j=1}^{N}\exp(r_{\theta}(x,y_{j}))}$ (24)
Given a human-provided listwise preference distribution $p^{*}_{i}$, we minimize the cross-entropy between $p^{*}_{i}$ and $P_{\theta}$:
$\mathcal{L}^{\text{DPO}}(\theta)=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\exp(r_{\theta}(x,y_{i}))}{\sum_{j=1}^{N}\exp(r_{\theta}(x,y_{j}))}\right)\right]$ (25)
Substituting the reward expression and simplifying:
$\mathcal{L}^{\text{DPO}}(\theta)&=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}Z(x)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}Z(x)^{\beta}}\right)\right]\\ &=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}\right)\right]\end{aligned}$
This becomes the canonical DPO loss. To extend this to multiple preference dimensions, we assume there are $m$ reward components indexed by $k$, each with its own preference distribution $p_{i}^{*(k)}$:
$\mathcal{L}^{(k)}(\theta)=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}\right)\right]$ (26)
We then define a user-controllable or sampled weight vector $\lambda\in\Delta^{m}$ (the $m$-dimensional probability simplex: $\lambda_{k}\geq 0$, $\sum_{k}\lambda_{k}=1$), and define the aggregated loss:
$\mathcal{L}^{\lambda}(\theta)=\sum_{k=1}^{m}\lambda_{k}\cdot\mathcal{L}^{(k)}(\theta)$ (27)
To enable generalization over arbitrary preferences, we optimize the expected loss over $\lambda\sim\text{Unif}(\Delta^{m})$:
$\mathcal{L}_{\text{final}}(\theta)=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\mathcal{L}^{\lambda}(\theta)\right]=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\sum_{k=1}^{m}\lambda_{k}\cdot\mathcal{L}^{(k)}(\theta)\right]$ (28)
A.2 From Final Objective to Lambda-DPO Form
To convert Eq. (25) to a form that operates on distributions directly, we change the order of summation. Recall:
$\mathcal{L}^{(k)}(\theta)=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (29)
So the aggregated expected loss becomes:
$\mathcal{L}_{\text{final}}(\theta)$ $=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\sum_{k=1}^{m}\lambda_{k}\cdot\mathbb{E}_{(x,\{y_{i}\})}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\log P_{\theta}(y_{i}|x)\right]\right]$ (30)
$=\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[-\sum_{i=1}^{N}\sum_{k=1}^{m}\lambda_{k}p_{i}^{*(k)}\log P_{\theta}(y_{i}|x)\right]$ (31)
We now define the aggregated distribution over outputs:
$p^{\lambda}(y_{i}|x)=\sum_{k=1}^{m}\lambda_{k}p^{*(k)}(y_{i}|x)$ (32)
So the objective becomes:
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (33)
Thus, by switching the summation order (i.e., aggregating over preference weights before computing loss), we arrive at an equivalent but more compact objective that is easier to optimize and sample in practice.
A.3 Gradient of Lambda-DPO Loss
We compute the gradient of $\mathcal{L}_{\lambda\text{-DPO}}$:
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}$ $=-\mathbb{E}_{x,\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\nabla_{\theta}\log P_{\theta}(y_{i}\mid x)\right]$ (34)
$=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$ (35)
where we use:
$\nabla_{\theta}P_{\theta}(y_{i}|x)=P_{\theta}(y_{i}|x)\cdot\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)$ (36)
This completes the derivation of the Lambda-weighted DPO objective and its gradient form.
Appendix B B Implementation of _compute_lambda_dpo_loss
This function executes the following procedure:
B.1 Input Tensors and Chunked Computation
  1. 1.
    Load the batch tensors:
    input_ids = inputs["input_ids"]
    attention_mask = inputs["attention_mask"]
    labels = inputs["labels"]
    pi_target = inputs["pi_target"]
  2. 2.
    Assume the batch contains 4 preference dimensions and 4 candidate responses (16 sequences in total). Reshape the tensors so that each preference dimension shares the same 4 unique responses:
    num_pref, num_resp = 4, 4
    B = input_ids.size(0) // (num_pref * num_resp)
    seq_len = input_ids.size(1)
    ids_view = input_ids.view(B, num_pref * num_resp, seq_len)[:, :num_resp, :].reshape(B * num_resp, seq_len)
    mask_view = attention_mask.view(B, num_pref * num_resp, seq_len)[:, :num_resp, :].reshape(B * num_resp, seq_len)
    labels_view = labels.view(B, num_pref * num_resp, seq_len)[:, :num_resp, :].reshape(B * num_resp, seq_len)
  3. 3.
    Choose a chunk size (by default equal to ids_view.size(0)) to compute log-probabilities in smaller batches. Optionally compute reference model probabilities:
    chunk_size = self.lambda_dpo_chunk_size or ids_view.size(0)
    seq_log_probs_list = []
    ref_seq_log_probs_list = []
    for start in range(0, ids_view.size(0), chunk_size):
    end = start + chunk_size
    chunk_ids = ids_view[start:end]
    chunk_mask = mask_view[start:end]
    chunk_labels = labels_view[start:end]
    logits = model(input_ids=chunk_ids, attention_mask=chunk_mask).logits
    seq_lp, valid_len = get_batch_logps(logits, chunk_labels, self.label_pad_token_id)
    seq_log_probs_list.append(seq_lp / valid_len)
    with torch.no_grad():
    if self.finetuning_args.use_ref_model and self.ref_model is not None:
    ref_logits = self.ref_model(input_ids=chunk_ids, attention_mask=chunk_mask).logits
    ref_lp, ref_len = get_batch_logps(ref_logits, chunk_labels, self.label_pad_token_id)
    ref_lp = ref_lp / ref_len
    else:
    ref_lp = torch.zeros_like(seq_lp)
    ref_seq_log_probs_list.append(ref_lp)
  4. 4.
    Concatenate the chunk results and share each set of four log-probabilities across the preference dimensions:
    seq_log_probs = torch.cat(seq_log_probs_list, dim=0)
    ref_seq_log_probs = torch.cat(ref_seq_log_probs_list, dim=0)
    seq_log_probs = seq_log_probs.view(B, num_resp).unsqueeze(1).expand(B, num_pref, num_resp).reshape(-1)
    ref_seq_log_probs = ref_seq_log_probs.view(B, num_resp).unsqueeze(1).expand(B, num_pref, num_resp).reshape(-1)
B.2 Reference Model and Relative Reward
  1. 5.
    Compute the relative reward and reshape back to $(B,4,4)$:
    r_theta = self.beta * (seq_log_probs - ref_seq_log_probs)
    r_theta = r_theta.view(B, num_pref, num_resp)
    pi_target = pi_target.view(B, num_pref, num_resp)
B.3 Listwise Loss by Preference Dimension
  1. 6.
    For each preference axis $i\in\{1,2,3,4\}$, compute the listwise loss:
    listwise_losses = []
    for i in range(num_pref):
    r = r_theta[:, i, :]
    p = pi_target[:, i, :] + 1e-10
    log_softmaxed = r - torch.logsumexp(r, dim=1, keepdim=True)
    loss = -(p * log_softmaxed).sum(dim=1).mean()
    listwise_losses.append(loss)
B.4 Lambda Weights and Final Loss
  1. 7.
    The current implementation equally weights the preference dimensions (no random sampling of $\lambda$):
    num_losses = len(listwise_losses)
    weights = torch.ones(num_losses, device=input_ids.device)
    weights /= weights.sum()
    final_loss = sum(w * l for w, l in zip(weights, listwise_losses))
    if return_outputs:
    return final_loss, {"loss_components": listwise_losses}
    return final_loss
Appendix C: Humorous Overview of the Added Classes
C.1 UltrafeedbackDatasetConverter
Input
A single UltraFeedback record. It contains:
  • a user prompt (instruction)
  • multiple completions, each with four ratings: helpfulness, honesty, instruction_following, and truthfulness.
Output
A dictionary with the prompt and, for every dimension, the raw scores plus a preference vector.
Example
Input:
instruction = "Explain photosynthesis."
completions = [
{"response": "...", "annotations": {"helpfulness": {"Rating": 4}}},
{"response": "...", "annotations": {"helpfulness": {"Rating": 2}}},
{"response": "...", "annotations": {"helpfulness": {"Rating": 5}}},
{"response": "...", "annotations": {"helpfulness": {"Rating": 3}}}
Output (simplified):
"_prompt": [{"role": "user", "content": "Explain photosynthesis."}],
"helpfulness": [
{"model": "?", "content": "...", "score": 4},
"_pi_target": {"helpfulness": [0.24, 0.09, 0.45, 0.22]}
C.2 ListwiseDatasetProcessor
Input
A dictionary produced by the converter that contains _prompt, the lists of responses for each dimension, and _pi_target.
Output
A flattened structure ready for training: for every sample there are 16 sequences of token IDs (one per response per dimension), matching labels and attention masks, and the list of probabilities.
Example
Input:
"_prompt": ...,
"helpfulness": [
{"content": "A"}, {"content": "B"}, {"content": "C"}, {"content": "D"}
"_pi_target": {"helpfulness": [0.24, 0.09, 0.45, 0.22], ...}
Output (illustrative shapes):
"input_ids": [seq1, seq2, ..., seq16], # each seq corresponds to a prompt+response
"labels": [...], # same layout as input_ids
"attention_mask": [...], # masks per sequence
"pi_target": [0.24, 0.09, 0.45, 0.22, ...] # flattened list of probabilities
C.3 ListwiseDataCollatorWithPadding
Input
A list of processed samples where each sample holds 16 sequences of token IDs and the flattened probability list.
Output
A padded batch containing tensors input_ids, attention_mask, labels, and an accompanying tensor of preference values aligned with the responses.
Example
Input:
# List of samples where each sample holds 16 sequences and
# a corresponding 16-element probability list
[example1, example2, ..., example_B] # B = batch size
Output:
"input_ids": Tensor of shape [B * 16, L],
"attention_mask": Tensor of shape [B * 16, L],
"labels": Tensor of shape [B * 16, L],
"pi_target": Tensor of shape [B * 16]
These three classes transform UltraFeedback data into tensors suitable for Lambda-weighted DPO training, while preserving the four-response structure and converting ratings into softmax preference vectors.
This function performs sequence probability calculation, reference model comparison, listwise cross-entropy over preference dimensions, and $\lambda$-weighted aggregation. It fully realizes the lambda-DPO loss objective described in the paper’s methodology section.
Appendix C D Implementation of _compute_ultra_dpo_loss
This routine computes the DPO loss when training on the UltraFeedback dataset, where each example contains four candidate responses with human preference scores. It proceeds as follows:
D.1 Input Aggregation
  1. 1.
    Load the tensors input_ids, attention_mask, labels, and pi_target.
  2. 2.
[…]
  3. 3.
    Apply uniform weights across the four preference dimensions to obtain an aggregated distribution for each example:
    $\text{aggregated}=\sum_{i=1}^{4}\tfrac{1}{4}\,\pi^{*(i)}.$
D.2 Selecting Best and Worst Responses
  1. 4.
    Determine the best response index via $\arg\max$ over aggregated and the worst index via $\arg\min$.
  2. 5.
    Gather the corresponding best and worst sequences from the original batch for further processing.
D.3 Log-Probability Computation
  1. 6.
    Pass the selected sequences through the model, compute token log-probabilities, and average over tokens to obtain sequence log-probabilities.
  2. 7.
    If a reference model is provided, compute its log-probabilities in the same way; otherwise use zero tensors as the reference scores.
D.4 Final DPO Loss
  1. 8.
    Split the policy and reference log-probabilities into best and worst halves.
  2. 9.
    Invoke compute_preference_loss on these values to obtain per-example DPO losses, then average across the batch:
    $\mathcal{L}_{\mathrm{DPO}}=\frac{1}{B}\sum_{b=1}^{B}\mathrm{loss}(y^{\mathrm{best}}_{b},\,y^{\mathrm{worst}}_{b}).$
Appendix D E Accuracy Comparisons on UltraFeedback Benchmarks with 1B model
We provide benchmark accuracy comparisons between DPO and $\lambda$-DPO across multiple language models on the UltraFeedback dataset. Performance is reported across six tasks: mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k.
E.1 Qwen-1.8 (Base)
Figure 5 shows that $\lambda$-DPO achieves modest improvements over standard DPO on most tasks using the Qwen-1.8 base model.
Figure 5: Accuracy comparison between DPO and $\lambda$-DPO on Qwen-1.8 (base).
E.2 Qwen-1.8 Chat
Figure 6 presents results on Qwen-1.8 Chat, where $\lambda$-DPO consistently improves accuracy, especially on reasoning-heavy benchmarks like arc and hellaswag.
Figure 6: Accuracy comparison between DPO and $\lambda$-DPO on Qwen-1.8 Chat.
E.3 LLaMA 3.2 1B (Base)
As shown in Figure 7, $\lambda$-DPO outperforms DPO slightly on LLaMA 3.2 1B base across four out of six tasks, demonstrating generalization beyond the Qwen series.
Figure 7: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B base model.
E.4 LLaMA 3.2 1B (Instruct)
Figure 8 illustrates results for the Instruct-tuned variant of LLaMA 3.2 1B. The performance gap between $\lambda$-DPO and DPO is generally preserved, indicating robustness to instruction tuning.
Figure 8: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B Instruct.
E.5 Unimodal Ablation
We visualize the performance of four unimodal $\lambda$-DPO models, each trained using a one-hot preference vector that activates only one of the four human preference dimensions: helpfulness, honesty, coherence, and fluency. Accuracy is reported across six standard benchmarks: mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k.
Figure 9: Accuracy comparison of four unimodal $\lambda$-DPO models, each trained with a single active preference dimension. All models perform similarly across benchmarks, with slight advantages for helpfulness and honesty on average. This suggests that no single preference dominates, and multi-dimensional interpolation may improve overall alignment.