### v5

# Multi-Preference Lambda-weighted Listwise DPO for Small-Scale Model Alignment

Abstract
Large language models (LLMs) demonstrate strong generalization across a wide range of language tasks, but often generate outputs that misalign with human preferences. Reinforcement Learning from Human Feedback (RLHF) addresses this by optimizing models toward human preferences using a learned reward function and reinforcement learning, yielding improved alignment but suffering from high computational cost and instability. Direct Preference Optimization (DPO) simplifies the process by treating alignment as a classification task over binary preference pairs, reducing training overhead while achieving competitive performance. However, it assumes fixed, single-dimensional preferences and only supports pairwise supervision. To address these limitations, we propose Multi-Preference Lambda-weighted Listwise DPO, which allows the model to learn from more detailed human feedback and flexibly balance multiple goals such as helpfulness, honesty, and fluency. Our method models full-ranked preference distributions rather than binary comparisons, enabling more informative learning signals. The $\lambda$ vector controls the relative importance of different alignment goals, allowing the model to generalize across diverse human objectives. During inference, $\lambda$ can be adjusted without retraining, providing controllable alignment behavior for downstream use. We also introduce a learned scheduler that dynamically samples performant $\lambda$ configurations to improve robustness. Notably, our method requires only 20GB of GPU memory for training, making it suitable for compute-constrained settings such as academic labs, educational tools, or on-device assistants. Experiments on 1B–2B scale models show that our method consistently outperforms standard DPO on alignment benchmarks while enabling efficient, controllable, and fine-grained adaptation suitable for real-world deployment.
Introduction
Large language models (LLMs) such as GPT (Brown et al. 2020) and LLaMA (Touvron and et al. 2023) have emerged as powerful backbones in natural language processing, demonstrating impressive capabilities across tasks such as question answering, summarization, and instruction following. However, despite their success, LLMs often produce outputs that do not align well with human expectations and values.
This misalignment frequently manifests as responses that are verbose, inaccurate, or unhelpful. It stems from their pretraining objective: next-token prediction, which emphasizes statistical co-occurrence over communicative intent, and inherently lacks exposure to explicit human feedback (Ouyang et al. 2022).
To address this, a line of alignment techniques has been developed to incorporate preference supervision. Among them, Reinforcement Learning from Human Feedback (RLHF) has become a standard paradigm (Christiano et al. 2017; Ouyang et al. 2022; Bai et al. 2022). By training a reward model from human comparisons and optimizing it via reinforcement learning (e.g., PPO (Schulman et al. 2017)), RLHF enables models to better match human values. However, the reliance on expensive reward modeling, unstable optimization dynamics, and high computational cost make RLHF difficult to scale and deploy reliably (Ziegler et al. 2019).
As a lightweight alternative, Direct Preference Optimization (DPO) has recently been proposed (Rafailov et al. 2023). DPO reframes alignment as a classification problem over pairwise preferences, directly optimizing model outputs to prefer preferred responses over less preferred ones. Without requiring an explicit reward model or reinforcement loop, DPO achieves strong alignment with reduced training overhead. Nevertheless, existing DPO methods typically assume a single, fixed alignment objective and binary supervision. This limits their applicability in practical scenarios, where alignment is inherently multi-dimensional—balancing objectives such as helpfulness, factuality, honesty, and fluency (Bai et al. 2022; Wu et al. 2021).
To accommodate this diversity, several controllable generation strategies have been explored, such as prompt tuning (Liu and et al. 2021), control tokens (Keskar et al. 2019), and multi-reward reinforcement learning (Lee et al. 2023). However, these approaches often require retraining under each new preference configuration, introduce architectural complexity, or generalize poorly in unseen alignment regimes.
In this work, we propose Multi-Preference Lambda-weighted Listwise DPO, a unified framework that extends the standard DPO formulation along two critical dimensions: supervision granularity and preference flexibility. Concretely, we generalize pairwise supervision to listwise preference modeling, enabling the model to learn from multiple ranked completions per prompt. This richer supervision signal allows the model to reason over global ranking structures, rather than relying on isolated binary comparisons. As a result, it improves gradient quality and better captures nuanced distinctions between high-quality and suboptimal responses. Simultaneously, we introduce a lambda-weighted composition mechanism, in which a preference vector $\lambda\in\Delta^{m}$ governs the relative importance of each alignment axis. This design allows a single model to flexibly interpolate between multiple objectives without retraining.
To further enhance adaptability, we develop a learned scheduling strategy that dynamically adjusts the sampling distribution over $\lambda$ vectors based on their empirical performance.
Empirical results on two strong 1B–2B scale LLMs (LLaMA and Qwen) across six alignment benchmarks demonstrate that our approach consistently outperforms standard DPO. Furthermore, by adjusting $\lambda$ during inference, our method supports personalized alignment without modifying model weights, making it highly practical for real-world deployment in education, conversational agents, and lightweight on-device applications.
Related Work
RLHF and DPO.
Reinforcement learning from human feedback (RLHF) (Ouyang et al. 2022; Bai et al. 2022) has become a widely used paradigm for aligning large language models (LLMs) with human preferences. It typically involves training a reward model using pairwise human comparisons and then optimizing the base language model using reinforcement learning algorithms such as Proximal Policy Optimization (PPO) (Schulman et al. 2017). This two-stage approach has demonstrated strong alignment performance, notably in systems like InstructGPT (Ouyang et al. 2022) and Claude (Bai et al. 2022). However, RLHF suffers from several limitations: reward models can be noisy or biased, PPO introduces high variance and instability during training, and the overall pipeline is computationally intensive and hard to scale (Ziegler et al. 2019; Christiano et al. 2017).
To reduce this complexity, direct preference optimization (DPO) (Rafailov et al. 2023) reframes preference alignment as a supervised classification problem. Instead of learning a reward model, DPO directly optimizes the policy to prefer preferred responses using a KL-regularized binary cross-entropy loss. This improves both training stability and implementation simplicity. Variants such as LLaMA-DPO (Liu et al. 2023) and Quark (Yao et al. 2023) have extended this framework with architectural tuning and reward quantization, showing strong performance in single-objective settings. However, most existing work in this area assumes fixed binary preferences and lacks the flexibility to accommodate multiple alignment goals or dynamically varying human objectives.
Listwise Preference Learning.
While pairwise preference modeling is common, listwise preference learning has been shown to better capture global ranking information (Burges et al. 2005; Xia et al. 2008). Instead of treating comparisons independently, listwise approaches model the entire ranking over a set of candidates, providing richer supervision and reducing sample complexity. In the context of language model alignment, UltraFeedback (Du et al. 2023) provides large-scale human feedback data with complete ranking annotations over multiple completions. This enables more expressive and fine-grained alignment signals beyond binary feedback.
Recent works such as GLOW (Xu et al. 2023) leverage such ranked data to train reward models using listwise objectives. These models are then used in RLHF-style pipelines to align policies. However, they still rely on intermediate reward learning and reinforcement optimization. In contrast, our method directly optimizes the policy using listwise cross-entropy loss, bypassing the need for explicit reward modeling. This not only simplifies the training pipeline but also provides more stable gradients and better captures the structure of human preferences.
Multi-Objective and Controllable Alignment.
Human alignment is inherently multi-faceted, involving trade-offs among objectives such as helpfulness, honesty, informativeness, and harmlessness. Prior work has explored multi-objective reinforcement learning frameworks (Korbak et al. 2023; Lee et al. 2023) that combine multiple reward signals through scalarization or decomposition. However, these methods typically require retraining or manual tuning of trade-off weights for each new configuration. Similarly, reward interpolation techniques (Wu et al. 2021) attempt to blend preferences using convex combinations of pretrained reward models, but they still involve separate optimization for each alignment setting.
To enable better generalization, our method integrates a weight vector $\lambda$ into the training process, allowing the model to internalize a continuous spectrum of preference trade-offs. This is conceptually related to universal value function approximators (Schaul et al. 2015), which generalize value functions across goals or reward vectors. In the alignment context, we treat $\lambda$ as a parameter controlling the target alignment behavior, enabling downstream control without model retraining.
Alternative strategies for controllable generation include prompt-based steering (Brown et al. 2020; Wei et al. 2022), control tokens (Keskar et al. 2019), and constrained decoding techniques (Lu et al. 2022). While effective in certain settings, these methods often struggle with consistency, lack semantic grounding in human-aligned dimensions, or require careful hand-engineering. Our approach differs in that it integrates multi-dimensional alignment directly into the learning objective, enabling fine-grained, interpretable control through $\lambda$ without modifying prompts, decoding strategies, or model architecture.
Preliminaries
We begin by reviewing the foundation of Direct Preference Optimization (DPO) and its connection to KL-regularized reinforcement learning from human feedback (RLHF). Our setting assumes the existence of a base language model $\pi_{\text{ref}}$, a supervised fine-tuned (SFT) model, and a collection of ranked or pairwise human preferences. We aim to train a target policy $\pi_{\theta}$ that better aligns with human preferences while remaining close to $\pi_{\text{ref}}$.
Pairwise Preference Modeling.
Let $x$ denote the prompt, and $y_{w}\succ y_{l}$ represent a human preference between a pair of completions: $y_{w}$ (the preferred answer) and $y_{l}$ (the less preferred one). These preferences are sampled from an underlying distribution $p^{*}(y_{w},y_{l}|x)$, which is typically assumed to be induced by a latent reward function $r(x,y)$. A commonly adopted model is the Bradley-Terry (BT) model (Bradley and Terry 1952):
$p^{*}(y_{w}\succ y_{l}\mid x)=\frac{\exp(r(x,y_{w}))}{\exp(r(x,y_{w}))+\exp(r(x,y_{l}))}$ (1)
Given a dataset of comparisons $\mathcal{D}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$, we can learn the reward function $r(x,y)$ by minimizing the negative log-likelihood under the BT model:
$\mathcal{L}_{\text{BT}}(r)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma(r(x,y_{w})-r(x,y_{l}))\right]$ (2)
where $\sigma(z)=1/(1+\exp(-z))$ is the logistic sigmoid.
KL-Regularized RL Objective.
Once a reward model is trained, the standard RLHF approach fine-tunes the policy $\pi_{\theta}$ to maximize expected rewards while constraining deviation from the reference policy $\pi_{\text{ref}}$ via KL regularization:
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}\left[r(x,y)-\beta\cdot\mathrm{KL}\left(\pi_{\theta}(y|x)\|\pi_{\text{ref}}(y|x)\right)\right]$ (3)
where $\beta>0$ controls the strength of the KL constraint.
This objective encourages the policy to produce high-reward outputs while staying close to $\pi_{\text{ref}}$. However, in practice, it requires sampling, reward evaluation, and reinforcement learning algorithms such as PPO (Schulman et al. 2017) to optimize.
Reparameterization and DPO.
DPO sidesteps explicit reward modeling and reinforcement learning by recognizing that the optimal policy $\pi_{r}$ under Equation 3 can be written as:
$\pi_{r}(y|x)\propto\pi_{\text{ref}}(y|x)\cdot\exp\left(\frac{1}{\beta}r(x,y)\right)$ (4)
Using Equation 4, we reparameterize the reward difference into the log-ratio of policies:
$\beta(r(x,y_{w})-r(x,y_{l}))=\log\frac{\pi_{r}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{r}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$ (5)
Replacing $\pi_{r}$ with $\pi_{\theta}$ yields the DPO loss:
$-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$ (6)
This loss directly optimizes the policy parameters $\theta$ using preference comparisons, without sampling or reinforcement learning. It forms the foundation upon which we build our multi-preference and listwise extensions in later sections.
Figure 1: Overview of the $\lambda$-DPO architecture. The system integrates listwise preference modeling and a controllable $\lambda$ vector to align model outputs with multi-dimensional human preferences.
Multi-Preference Lambda-weighted Listwise DPO
Motivated by the limitations of pairwise DPO under a single fixed preference distribution, we propose two orthogonal enhancements: listwise supervision and simplex-weighted optimization. First, instead of relying on binary comparisons, we adopt listwise learning to utilize full human rankings over multiple outputs. This provides stronger supervision signals and captures richer preference structure. Second, we generalize DPO to support multi-objective alignment by introducing a simplex-weighted aggregation over multiple reward components. The resulting $\lambda$-DPO formulation enables flexible interpolation across alignment dimensions such as helpfulness, harmlessness, and fluency. Together, these extensions allow controllable and scalable alignment with more nuanced human feedback.
Listwise Preference Modeling.
Let $\{y_{1},y_{2},\dots,y_{N}\}$ be a set of candidate completions for a prompt $x$, and let $p^{*}(y_{i}|x)$ be a listwise human preference distribution over these candidates. Analogous to the pairwise DPO reparameterization, we define the listwise policy distribution under $\pi_{\theta}$ as:
$P_{\theta}(y_{i}|x)=\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}$ (7)
which defines a softmax over reweighted log-ratios, consistent with KL-derived optimal policies.
We then define the listwise DPO objective as a cross-entropy loss between human preferences $p^{*}(y_{i}|x)$ and model predictions:
$\mathcal{L}_{\text{listwise}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[\sum_{i=1}^{N}p^{*}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right]$ (8)
Lambda-weighted Preference Composition.
To enable multi-objective control, we assume access to $m$ distinct preference dimensions, each defined by a distribution $p^{*(k)}(y_{i}|x)$ for objective $k\in\{1,\dots,m\}$. For each prompt, we sample or specify a weight vector $\lambda\in\Delta^{m}$ (the probability simplex), and define the aggregated preference distribution:
$p^{\lambda}(y_{i}|x)=\sum_{k=1}^{m}\lambda_{k}\cdot p^{*(k)}(y_{i}|x)$ (9)
This allows us to interpolate among various alignment goals without retraining.
The full Lambda-weighted Listwise DPO objective becomes:
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right]$ (10)
This expectation can be approximated by sampling $\lambda\sim\text{Unif}(\Delta^{m})$ or conditioning on user-specified preferences.
What does $\lambda$-DPO update do?
To understand the optimization dynamics, consider the gradient of the loss:
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}|x)-P_{\theta}(y_{i}|x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\right]$ (11)
This gradient updates the model to increase likelihood of responses aligned with weighted human preferences, while decreasing likelihood of responses overrepresented in the reference policy. Importantly, when $\lambda$ is fixed (e.g., $\lambda=(1,0,0)$), the objective reduces to a single-objective listwise DPO loss.
Lambda-DPO Pipeline.
The training process proceeds as follows:
  1. 1.
    For each prompt $x$, obtain $N$ candidate completions $\{y_{1},\dots,y_{N}\}\sim\pi_{\text{ref}}(\cdot|x)$.
  2. 2.
    For each alignment dimension $k$, elicit a human listwise preference distribution $p^{*(k)}(y_{i}|x)$.
  3. 3.
    Sample or define $\lambda\in\Delta^{m}$, and compute the mixture preference $p^{\lambda}$.
  4. 4.
    Optimize $\pi_{\theta}$ using the listwise cross-entropy loss between $p^{\lambda}$ and $P_{\theta}$.
In practice, we can reuse datasets like UltraFeedback (Du et al. 2023), where multiple preference criteria are annotated. During inference, changing $\lambda$ enables personalized alignment without modifying model parameters.
Theoretical Analysis of Multi-Preference Lambda-weighted Listwise DPO
While Direct Preference Optimization (DPO) offers a principled and scalable approach to aligning language models with human feedback, it is fundamentally limited by its reliance on pairwise supervision. In this section, we provide a detailed theoretical analysis of these limitations and motivate our proposed extensions—namely, listwise supervision and preference mixing via simplex weighting—to overcome them in multi-objective alignment settings.
Pairwise DPO and Its Limitations.
Given a prompt $x$ and a pair of sampled completions $y_{w}$ (preferred) and $y_{l}$ (dispreferred), DPO defines a binary target distribution over the pair as:
$p^{*}(y_{i}\mid x)=\begin{cases}1&\text{if }y_{i}=y_{w}\\ 0&\text{if }y_{i}=y_{l}\end{cases}$ (12)
This construction is simple and efficient, allowing the loss to be computed from binary human preferences. However, it treats the preference signal as a hard decision, discarding nuanced feedback such as “somewhat better” or “equally good.” Furthermore, the normalization over only two samples induces high-variance gradients, as the updates depend on stochastic comparisons between limited completions. Since only the relative preference between $y_{w}$ and $y_{l}$ is modeled, all other completions (which might be equally or more suitable) are ignored. This can lead to suboptimal generalization, especially when the number of plausible high-quality completions is large.
Listwise Distribution and Multi-Preference Supervision.
To address the expressiveness gap, we generalize DPO to operate over a list of completions $\{y_{1},\dots,y_{N}\}$ using listwise supervision. Specifically, given a reward model $r(x,y_{i})$ that quantifies the quality of each completion, we define a normalized target distribution over the list using softmax:
$p^{*}(y_{i}\mid x)=\frac{\exp(r(x,y_{i}))}{\sum_{j=1}^{N}\exp(r(x,y_{j}))}$ (13)
This formulation assigns probability mass proportionally to reward values, providing smoother gradients and improving credit assignment. Unlike the binary setting, listwise DPO makes use of the full ranking structure, offering a richer training signal and faster convergence.
Moreover, human preferences in practical applications are inherently multi-dimensional—comprising objectives like helpfulness, safety, factuality, politeness, and conciseness. Suppose we define $m$ distinct reward models, each representing one objective. We can construct a family of listwise distributions $\{p^{*(k)}(y_{i}|x)\}_{k=1}^{m}$, each aligned with a single criterion. A convex combination of these objectives is parameterized by a preference vector $\lambda\in\Delta^{m}$, forming a new composite distribution:
$p^{\lambda}(y_{i}\mid x)=\sum_{k=1}^{m}\lambda_{k}\cdot p^{*(k)}(y_{i}\mid x)$ (14)
This formulation enables controllable alignment by adjusting $\lambda$, allowing practitioners to emphasize certain values over others (e.g., prioritize safety over informativeness). The model is trained to match this target distribution via cross-entropy:
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (15)
Its gradient reveals the direction of improvement, given by $\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=$
$-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$ (16)
which encourages the model’s policy distribution to match the aggregated preference signal from multiple human value dimensions.
Summary.
In summary, Multi-Preference Lambda-weighted Listwise DPO offers a theoretically grounded generalization of the original DPO framework. First, it extends binary supervision to listwise distributions, providing smoother and more informative gradient updates. Second, it integrates diverse human objectives through a simplex-weighted mixture of reward-aligned targets, enabling dynamic and tunable preference alignment. Finally, it reduces sample inefficiency and training instability by leveraging full output rankings and avoiding overfitting to binary labels, thus making it well-suited for real-world alignment tasks with conflicting value axes.
Comparative Experiment
We compare standard DPO and our proposed $\lambda$-DPO under identical training and evaluation setups on two models: LLaMA 3.2-1B Instruct and Qwen 1.8B Chat. Both models are first supervised fine-tuned (SFT) on UltraChat-200k, then aligned using the UltraFeedback listwise dataset.
Training Details.
We use LoRA (rank 8, all-layer target) with a learning rate of $5\times 10^{-6}$, batch size 1, gradient accumulation steps 8, cosine learning rate schedule, and 5 epochs. The KL penalty $\beta$ is set to 0.1. $\lambda$-DPO uses the lambda_dpo loss with chunk size 1.
Main Results
Table 1 reports the accuracy of DPO and $\lambda$-DPO on six representative benchmarks for both LLaMA 3.2-1B Instruct and Qwen 1.8B Chat. On LLaMA, $\lambda$-DPO consistently improves performance across all tasks, with notable gains such as +1.1% on ARC, +3.3% on GSM8K, and +3.6% on TruthfulQA. The improvement on WSC273 is also evident (+1.8%). For Qwen 1.8B Chat, $\lambda$-DPO outperforms DPO on all benchmarks except TruthfulQA, where DPO achieves a slightly higher score. Notably, large gains are observed on GSM8K (+14.7%) and MMLU (+3.9%), demonstrating the generalization benefits of $\lambda$-DPO across different model families.
Model Method MMLU ARC HellaSwag TruthfulQA WSC273 GSM8K
LLaMA 3.2-1B Instruct ORPO 0.456 0.355 0.447 0.435 0.678 0.327
SIMPO 0.459 0.362 0.455 0.381 0.703 0.318
DPO 0.458 0.364 0.456 0.378 0.696 0.312
$\lambda$-DPO 0.460 0.375 0.459 0.414 0.714 0.345
Qwen 1.8B Chat ORPO 0.377 0.278 0.417 0.347 0.661 0.007
SIMPO 0.415 0.295 0.423 0.355 0.658 0.007
DPO 0.379 0.271 0.418 0.366 0.656 0.008
$\lambda$-DPO 0.418 0.284 0.428 0.312 0.667 0.155
Table 1: Accuracy comparison across six benchmarks. The first block compares ORPO, SIMPO, DPO, and $\lambda$-DPO on LLaMA 3.2-1B Instruct. The second block compares the same four methods on Qwen 1.8B Chat. Bold values indicate the best performance within each model group.
Visualization
To further illustrate these results, we provide bar charts comparing DPO and $\lambda$-DPO performance on the two models.
Figure 2: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B Instruct.
Figure 3: Accuracy comparison between DPO and $\lambda$-DPO on Qwen 1.8B Chat.
Performance Analysis and Discussion
As shown in both Table 1 and Figures 23, $\lambda$-DPO consistently improves performance across most tasks. The method yields notable improvements on ARC, GSM8K, and MMLU, while maintaining parity on Winogrande. TruthfulQA is the only benchmark where DPO slightly outperforms $\lambda$-DPO (in the Qwen model).
These results suggest that $\lambda$-DPO effectively captures human preference signals across diverse task types. Its listwise preference modeling contributes to better alignment with human-evaluated quality metrics such as factuality, helpfulness, and fluency.
Nonetheless, for certain adversarial or logic-heavy tasks like TruthfulQA, the benefits of listwise supervision may be less reliable. Integrating $\lambda$-DPO with structured supervision—such as chain-of-thought prompting (Wei et al. 2022)—could potentially mitigate this gap.
Conclusion of Comparative Analysis
In summary, $\lambda$-DPO brings two key improvements over standard DPO. First, by adopting listwise preference supervision, it leverages full ranking information instead of pairwise comparisons, resulting in richer and lower-variance gradient signals. Second, through the introduction of the $\lambda$ vector, it supports flexible interpolation across multiple alignment objectives (e.g., helpfulness, harmlessness, factuality), enabling controllable alignment behavior and improved generalization under diverse user preferences. These two enhancements jointly contribute to the observed performance gains across benchmarks. Future work may explore hybrid strategies that combine $\lambda$-DPO with structured reasoning techniques to further boost alignment quality on complex tasks.
Ablation Experiment
To further examine the effectiveness of different preference vector settings, we conduct an ablation experiment focusing on four unimodal $\lambda$-DPO models. Each model is trained using a fixed one-hot $\lambda$ vector, targeting only one of the four annotated dimensions: helpfulness, honesty, instruction-following, and truthfulness. This setup allows us to isolate the influence of each individual preference axis and compare their respective impacts across evaluation benchmarks.
Unimodal Comparison Across Benchmarks
We evaluate the four unimodal $\lambda$-DPO variants across six standard benchmarks—mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k. Figure 4 shows the accuracy distribution of each model across these tasks.
Figure 4: Accuracy Comparison of Unimodal $\lambda$-DPO Models with Fixed Preference Vectors.
All four models perform similarly across benchmarks. The honesty-focused model ($\lambda=[0,1,0,0]$) achieves the highest average accuracy, while the helpfulness-focused model ($\lambda=[1,0,0,0]$) excels on truthfulqa and performs well on gsm8k. Other preference models show only slight drops, with no significant performance degradation.
Preference Ranking and Vector Hypothesis
To better assess how each individual preference dimension contributes to generalization performance, we compute the average accuracy of the four unimodal $\lambda$-DPO models across six benchmark tasks:
  • $\lambda=[1,0,0,0]$: 0.4593
  • $\lambda=[0,1,0,0]$: 0.4598
  • $\lambda=[0,0,1,0]$: 0.4543
  • $\lambda=[0,0,0,1]$: 0.4548
  • $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$: 0.4612
The results indicate that all four unimodal preference dimensions contribute to model alignment, with helpfulness and honesty yielding slightly better generalization. However, the uniform interpolation $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ achieves the highest overall accuracy (0.4612), outperforming all unimodal variants.
This highlights the benefit of explicitly retaining the $\lambda$ vector for multi-objective supervision. Unlike standard DPO, which implicitly collapses preferences into a fixed form, $\lambda$-DPO provides fine-grained control over alignment trade-offs and leads to more robust generalization.
Discussion and Ablation Analysis
To further understand how individual preference dimensions influence alignment, we observe from Figure 4 that:
  • The $\lambda=[1,0,0,0]$ (helpfulness) model performs best on truthfulqa and achieves solid results on gsm8k.
  • The $\lambda=[0,1,0,0]$ (honesty) model leads slightly in overall average accuracy, especially strong on mmlu.
  • The $\lambda=[0,0,1,0]$ (instruction-following) and $\lambda=[0,0,0,1]$ (fluency) models show close and competitive performance, contrary to the notion that they are stylistically weaker.
These results indicate that optimizing for either content-driven or stylistic objectives in isolation yields comparable outcomes under our listwise framework. It further motivates the integration of multiple preference signals, especially for tasks requiring both correctness and user-aligned expressiveness.
Performance-Driven Lambda Scheduling via Polynomial Modeling
Performance-Driven Scheduler Design
We propose a performance-driven scheduler to improve multi-objective alignment by dynamically selecting preference vectors $\lambda$ based on their estimated effectiveness. Instead of fixing a single $\lambda$, the scheduler learns a distribution over $\lambda$ configurations that reflect observed performance.
The scheduler fits a polynomial regression model to past $(\lambda,\text{accuracy})$ pairs and uses it to score new candidates sampled from a Dirichlet prior over the simplex. A softmax over predicted scores then defines sampling probabilities, balancing exploitation of high-performing regions and exploration. This adaptive strategy promotes training diversity and improves robustness.
Empirical Comparison with Uniform Baseline
As shown in Figure 12, the learned scheduler achieves slightly better average accuracy (0.462 vs. 0.461) across six evaluation benchmarks when compared with a static baseline that uses the uniform configuration $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$. While the performance gain is modest, the scheduler enhances robustness and reduces overfitting by continuously adapting to observed training signals.
Importantly, this framework supports flexible preference adaptation and can be extended to personalized or task-specific settings without requiring retraining.
Figure 5: Comparison between learned $\lambda$ scheduler and uniform $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ across six alignment benchmarks. The learned scheduler achieves a slightly higher average accuracy and improves robustness by distributing preference sampling over higher-quality regions of the simplex.
Extension to LLaMA-8B
To assess the scalability of $\lambda$-DPO, we extend our experiments to the LLaMA-8B model. Across six standard benchmarks, $\lambda$-DPO yields an average relative improvement of approximately 0.3% over DPO, compared to a much larger average gain of 1.7% on the LLaMA-1B model.
Figure 6: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA-8B across six benchmarks.
As shown in Figure 6, $\lambda$-DPO achieves better alignment on the 1B model, with notably larger gains on challenging tasks such as GSM8K. This suggests that smaller models benefit more from listwise multi-preference supervision, likely because alignment signals have a stronger relative effect when model capacity is limited.
Moreover, training the 1B model only requires around 20GB of GPU memory, making it both efficient and accessible for academic labs, startups, or edge-focused applications. This low resource footprint enables rapid experimentation and deployment on commodity hardware, without sacrificing alignment quality. As such, $\lambda$-DPO is well-suited for scalable, real-world use cases where cost, speed, and controllability are critical.
Conclusion
In this paper, we propose Multi-Preference Lambda-weighted Listwise DPO ($\lambda$-DPO) to enable efficient, controllable, and fine-grained alignment of language models with diverse human preferences.
For robust and flexible alignment, we generalize pairwise supervision to listwise preference modeling, allowing the model to learn from ranked completions and capture nuanced preference structures. To support personalized and multi-objective alignment, we introduce a simplex-weighted $\lambda$ vector that dynamically balances objectives such as helpfulness, honesty, and fluency—without retraining.
To enhance generalization and robustness, we incorporate a learned scheduler that adaptively samples high-performing $\lambda$ configurations during training, improving diversity while mitigating modeling bias. Additionally, the method is designed to be resource-efficient, requiring only 20GB of GPU memory, making it practical for deployment in constrained scenarios such as on-device assistants and educational tools.
Experimental results on LLaMA 3.2-1B and Qwen-1.8B demonstrate that $\lambda$-DPO achieves consistent improvements over standard DPO, with average accuracy gains of +1.7% and +2.8%, respectively, across standard alignment benchmarks.
Acknowledgements
We thank the contributors of the original DPO framework for their foundational work, which inspired this extension. We also appreciate the open-source efforts of the Hugging Face team and the LLaMA-Factory project for enabling scalable preference-based fine-tuning.
This work was made possible through collaborative effort and self-supported compute resources by the authors: Yuhui Sun, Xiyao Wang, Zixi Li, Zhenlong Yuan and Jinman Zhao. We further thank colleagues and early readers for helpful feedback that improved the theoretical formulation and presentation.
Appendix A A Mathematical Derivations
A.1 Deriving the Final Training Objective
We begin by defining the base reward model used in DPO. Let $\pi_{\theta}(y|x)$ denote the current policy and $\pi_{\text{ref}}(y|x)$ a fixed reference policy. Given input $x$ and candidate outputs $\{y_{1},\ldots,y_{N}\}$, the reward is defined via KL divergence as:
$r_{\theta}(x,y)=\beta\log\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)+\beta\log Z(x)$ (17)
where $Z(x)$ is a normalizing constant ensuring scale-invariance:
$Z(x)=\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)$ (18)
The predicted distribution over completions is then:
$P_{\theta}(y_{i}|x)=\frac{\exp(r_{\theta}(x,y_{i}))}{\sum_{j=1}^{N}\exp(r_{\theta}(x,y_{j}))}$ (19)
Given a human-provided listwise preference distribution $p^{*}_{i}$, we minimize the cross-entropy between $p^{*}_{i}$ and $P_{\theta}$:
$\mathcal{L}^{\text{DPO}}(\theta)=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\exp(r_{\theta}(x,y_{i}))}{\sum_{j=1}^{N}\exp(r_{\theta}(x,y_{j}))}\right)\right]$ (20)
Substituting the reward expression and simplifying:
$\mathcal{L}^{\text{DPO}}(\theta)&=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}Z(x)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}Z(x)^{\beta}}\right)\right]\\ &=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}\right)\right]\end{aligned}$
This becomes the canonical DPO loss. To extend this to multiple preference dimensions, we assume there are $m$ reward components indexed by $k$, each with its own preference distribution $p_{i}^{*(k)}$:
$\mathcal{L}^{(k)}(\theta)=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}\right)\right]$ (21)
We then define a user-controllable or sampled weight vector $\lambda\in\Delta^{m}$ (the $m$-dimensional probability simplex: $\lambda_{k}\geq 0$, $\sum_{k}\lambda_{k}=1$), and define the aggregated loss:
$\mathcal{L}^{\lambda}(\theta)=\sum_{k=1}^{m}\lambda_{k}\cdot\mathcal{L}^{(k)}(\theta)$ (22)
To enable generalization over arbitrary preferences, we optimize the expected loss over $\lambda\sim\text{Unif}(\Delta^{m})$:
$\mathcal{L}_{\text{final}}(\theta)=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\mathcal{L}^{\lambda}(\theta)\right]=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\sum_{k=1}^{m}\lambda_{k}\cdot\mathcal{L}^{(k)}(\theta)\right]$ (23)
A.2 From Final Objective to Lambda-DPO Form
To convert Eq. (25) to a form that operates on distributions directly, we change the order of summation. Recall:
$\mathcal{L}^{(k)}(\theta)=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (24)
So the aggregated expected loss becomes:
$\mathcal{L}_{\text{final}}(\theta)$ $=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\sum_{k=1}^{m}\lambda_{k}\cdot\mathbb{E}_{(x,\{y_{i}\})}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\log P_{\theta}(y_{i}|x)\right]\right]$ (25)
$=\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[-\sum_{i=1}^{N}\sum_{k=1}^{m}\lambda_{k}p_{i}^{*(k)}\log P_{\theta}(y_{i}|x)\right]$ (26)
We now define the aggregated distribution over outputs:
$p^{\lambda}(y_{i}|x)=\sum_{k=1}^{m}\lambda_{k}p^{*(k)}(y_{i}|x)$ (27)
So the objective becomes:
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (28)
Thus, by switching the summation order (i.e., aggregating over preference weights before computing loss), we arrive at an equivalent but more compact objective that is easier to optimize and sample in practice.
A.3 Gradient of Lambda-DPO Loss
We compute the gradient of $\mathcal{L}_{\lambda\text{-DPO}}$:
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}$ $=-\mathbb{E}_{x,\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\nabla_{\theta}\log P_{\theta}(y_{i}\mid x)\right]$ (29)
$=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$
where we use:
$\nabla_{\theta}P_{\theta}(y_{i}|x)=P_{\theta}(y_{i}|x)\cdot\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)$ (30)
This completes the derivation of the Lambda-weighted DPO objective and its gradient form.
Appendix B B Experiment Setup
B.1 Experimental Objectives and Comparison Targets
The primary goal of our experiments is to evaluate the performance of the proposed $\lambda$-DPO algorithm compared to the standard DPO method. Specifically, we assess their effectiveness across different model architectures and training settings using public preference datasets and standard evaluation benchmarks.
We consider two levels of comparison:
Model comparison.
We experiment with two strong open-source language models—LLaMA and Qwen. These models differ in architecture, pretraining corpora, and tokenization schemes, providing a robust testbed for evaluating generalizability.
Method comparison.
We compare DPO and $\lambda$-DPO. For each method-model pair, we perform independent hyperparameter tuning (e.g., learning rate, batch size, $\beta$) to ensure comparisons are fair and reflect the best achievable performance under comparable resource constraints.
B.2 Training Settings
Inspired by the SimPO baseline, we adopt a unified training pipeline:
  • Finetuned-Base: The model is first SFT-trained on UltraChat-200k, then optimized on UltraFeedback via DPO or $\lambda$-DPO to reflect realistic alignment scenarios.
B.3 Evaluation Benchmarks and Metrics
We evaluate model performance using the lm-eval-harness (Gao et al. 2024) framework, which reports accuracy on diverse benchmarks covering reasoning, factual correctness, and commonsense understanding.
Specifically, we report average accuracy across the following six established datasets:
  • MMLU: 57 subjects, 15K questions assessing academic knowledge (Hendrycks et al. 2020).
  • ARC: 1.1K grade-school science questions requiring reasoning (Clark et al. 2018).
  • HellaSwag: 10K commonsense sentence completions with distractors (Zellers et al. 2019).
  • TruthfulQA: 817 questions testing factuality and hallucination (Lin, Hilton, and Evans 2021).
  • Winograd: 273 examples probing coreference and syntax (Levesque, Davis, and Morgenstern 2012).
  • GSM8K: 8.5K arithmetic reasoning problems (Cobbe et al. 2021).
B.4 Ablation: Impact of Lambda Weighting
To better understand the impact of the $\lambda$ vector in $\lambda$-DPO, we conduct ablation experiments focusing on how different preference weightings affect alignment performance.
Specifically, we compare two $\lambda$ schedule variants within the listwise $\lambda$-DPO framework:
  • Uniform $\lambda$: Equal weights across all dimensions ($\lambda=(1/m,\dots,1/m)$), simulating balanced multi-objective aggregation.
  • Unimodal $\lambda$: One-hot vectors (e.g., $(1,0,\dots)$) isolating single-dimension optimization without interpolation.
This comparison helps isolate the contribution of $\lambda$-based preference interpolation within the listwise setting.
Appendix C C Accuracy Comparisons on UltraFeedback Benchmarks with 1B model
We provide benchmark accuracy comparisons between DPO and $\lambda$-DPO across multiple language models on the UltraFeedback dataset. Performance is reported across six tasks: mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k.
C.1 Qwen-1.8 (Base)
Figure 7 shows that $\lambda$-DPO achieves modest improvements over standard DPO on most tasks using the Qwen-1.8 base model.
Figure 7: Accuracy comparison between DPO and $\lambda$-DPO on Qwen-1.8 (base).
C.2 Qwen-1.8 Chat
Figure 8 presents results on Qwen-1.8 Chat, where $\lambda$-DPO consistently improves accuracy, especially on reasoning-heavy benchmarks like arc and hellaswag.
Figure 8: Accuracy comparison between DPO and $\lambda$-DPO on Qwen-1.8 Chat.
C.3 LLaMA 3.2 1B (Base)
As shown in Figure 9, $\lambda$-DPO outperforms DPO slightly on LLaMA 3.2 1B base across four out of six tasks, demonstrating generalization beyond the Qwen series.
Figure 9: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B base model.
C.4 LLaMA 3.2 1B (Instruct)
Figure 10 illustrates results for the Instruct-tuned variant of LLaMA 3.2 1B. The performance gap between $\lambda$-DPO and DPO is generally preserved, indicating robustness to instruction tuning.
Figure 10: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B Instruct.
C.5 Unimodal Ablation
We visualize the performance of four unimodal $\lambda$-DPO models, each trained using a one-hot preference vector that activates only one of the four human preference dimensions: helpfulness, honesty, coherence, and fluency. Accuracy is reported across six standard benchmarks: mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k.
Figure 11: Accuracy comparison of four unimodal $\lambda$-DPO models, each trained with a single active preference dimension. All models perform similarly across benchmarks, with slight advantages for helpfulness and honesty on average. This suggests that no single preference dominates, and multi-dimensional interpolation may improve overall alignment.
Appendix D D Performance-Driven Lambda Scheduling via Polynomial Modeling
We introduce a general and flexible method for performance-driven lambda scheduling, which constructs a probability distribution over preference vectors $\lambda\in\Delta^{d}$ based on observed alignment performance. Each supervised training point $(\lambda^{(i)},y^{(i)})$ consists of a preference configuration $\lambda^{(i)}$ and a corresponding scalar performance metric $y^{(i)}$ (e.g., average benchmark accuracy). The scheduling process unfolds in four steps:
D.1 Polynomial Feature Mapping
We begin by mapping each preference vector $\lambda$ to a high-dimensional feature space using a polynomial embedding. Specifically, we define a degree-$p$ polynomial feature map $\phi_{p}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N}$ that enumerates all monomials up to degree $p$:
$\phi_{p}(\lambda)=\left[1,\lambda_{1},\dots,\lambda_{d},\lambda_{1}^{2},\lambda_{1}\lambda_{2},\dots,\lambda_{d}^{p}\right]^{\top},$ (31)
where the total number of features is given by:
$N=\binom{d+p}{p}.$ (32)
This expansion enables expressive yet tractable modeling of non-linear relationships between $\lambda$ vectors and alignment performance.
D.2 Polynomial Regression
Using the training set $\{(\lambda^{(i)},y^{(i)})\}_{i=1}^{n}$, we fit a least-squares regression model in the lifted feature space to learn the performance landscape:
$\min_{\mathbf{w}\in\mathbb{R}^{N}}\sum_{i=1}^{n}\left(\mathbf{w}^{\top}\phi_{p}(\lambda^{(i)})-y^{(i)}\right)^{2}.$ (33)
The resulting predictive function
$f(\lambda)=\mathbf{w}^{\top}\phi_{p}(\lambda)$ (34)
approximates the expected performance of any candidate $\lambda$ on downstream alignment tasks.
D.3 Grid Construction over the Simplex
To enable tractable sampling from the $\lambda$ space, we construct a discrete candidate set $\{\lambda^{(j)}\}_{j=1}^{k}\subset\Delta^{d}$ via one of the following methods:
  • Simplex discretization: generate all combinations with replacement and normalize;
  • Dirichlet sampling: draw $\lambda^{(j)}\sim\mathrm{Dir}(\boldsymbol{\alpha})$, where $\boldsymbol{\alpha}=\mathbf{1}$ yields a uniform prior over the simplex.
D.4 Softmax-Based Scheduling Distribution
We convert the predicted scores $s_{j}=f(\lambda^{(j)})$ into a probability distribution over the candidate $\lambda$ vectors using a softmax transformation:
$p(\lambda^{(j)})=\frac{\exp(\tau\cdot f(\lambda^{(j)}))}{\sum_{l=1}^{k}\exp(\tau\cdot f(\lambda^{(l)}))},$ (35)
where the temperature parameter $\tau>0$ controls the exploration–exploitation trade-off.
D.5 Empirical Instantiation with Five Training Points
To validate the polynomial scheduling framework, we use five observed training points—four unimodal $\lambda$ vectors and one uniform configuration—each paired with an empirical alignment accuracy:
$\lambda^{(1)}$ $=[1,0,0,0],\quad y^{(1)}=0.4593$
$\lambda^{(2)}$ $=[0,1,0,0],\quad y^{(2)}=0.4598$
$\lambda^{(3)}$ $=[0,0,1,0],\quad y^{(3)}=0.4543$
$\lambda^{(4)}$ $=[0,0,0,1],\quad y^{(4)}=0.4548$
$\lambda^{(5)}$ $=[0.25,0.25,0.25,0.25],\quad y^{(5)}=0.4612$
We set $p=2$, yielding $N=\binom{4+2}{2}=15$ features. Using gradient descent to minimize squared error, we obtain:
$f(\lambda)$ $=0.35+0.09\lambda_{1}+0.09\lambda_{2}+0.09\lambda_{3}+0.09\lambda_{4}$
$+0.02\lambda_{1}^{2}+0.04\lambda_{1}\lambda_{2}+0.03\lambda_{1}\lambda_{3}+0.02\lambda_{1}\lambda_{4}$
$+0.02\lambda_{2}^{2}+0.02\lambda_{2}\lambda_{3}+0.02\lambda_{2}\lambda_{4}$
$+0.01\lambda_{3}^{2}+0.00\lambda_{3}\lambda_{4}+0.01\lambda_{4}^{2}$
D.6 Lambda Sampling from Learned Scheduler
We sample $k=10$ $\lambda$ vectors from Dirichlet(1,1,1,1)$1$(1,1,1,1)( 1 , 1 , 1 , 1 ) and compute predicted scores $f(\lambda^{(j)})$, converting them via softmax ($\tau=100$):
$p(\lambda^{(j)})=\frac{\exp(\tau\cdot f(\lambda^{(j)}))}{\sum_{l=1}^{10}\exp(\tau\cdot f(\lambda^{(l)}))}$
$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $f(\lambda)$ $p(\lambda)$
0.212 0.334 0.245 0.209 0.452 0.095
0.126 0.237 0.131 0.507 0.450 0.075
0.542 0.079 0.256 0.123 0.455 0.117
0.233 0.721 0.020 0.025 0.459 0.185
0.004 0.334 0.281 0.381 0.449 0.067
0.320 0.069 0.513 0.099 0.451 0.082
0.236 0.155 0.358 0.251 0.452 0.093
0.141 0.110 0.701 0.048 0.454 0.110
0.070 0.198 0.388 0.344 0.451 0.081
0.090 0.139 0.306 0.465 0.450 0.095
Table 2: Sampled $\lambda$ vectors, their predicted scores $f(\lambda)$, and softmax sampling probabilities $p(\lambda)$ at temperature $\tau=100$.
D.7 Scheduler vs. Uniform Comparison
We compare our learned scheduler with a uniform $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ baseline using LLaMA 3.2B Instruct. The learned scheduler achieves slightly better average accuracy (0.462 vs. 0.461), and offers regularization via stochastic $\lambda$ sampling, enhancing robustness.
Figure 12: Comparison of learned scheduler vs. fixed uniform $\lambda$ across six alignment benchmarks. The learned scheduler yields improved accuracy and diversity.